### v2

# The $\alpha$–regression for compositional data: a unified framework for standard, spatially-lagged, spatial autoregressive and geographically-weighted regression models

Abstract
Compositional data—vectors of non-negative components summing to unity—frequently arise in scientific applications where covariates influence the relative proportions of components, yet traditional regression approaches ace challenges regarding the unit-sum constraint and zero values. This paper revisits the $\alpha$–regression framework, which uses a flexible power transformation parameterized by $\alpha$ to interpolate between raw data analysis and log-ratio methods, naturally handling zeros without imputation while allowing data-driven transformation selection. We formulate $\alpha$–regression as a non-linear least squares problem, provide efficient estimation via the Levenberg-Marquardt algorithm, and derive marginal effects for interpretation. The framework is extended to spatial settings through two models: the $\alpha$–spatially lagged X regression model, which incorporates spatial spillover effects via spatially lagged covariates with decomposition into direct and indirect effects, the $\alpha$–spatially autoregressive regression model and the geographically weighted $\alpha$–regression, which allows coefficients to vary spatially for capturing local relationships. Applications to two real data sets illustrate the performance of the models and showcase that spatial extensions capture the spatial dependence and improve the predictive performance.
Keywords: compositional data, $\alpha$–transformation, spatial regression
1 Introduction
Compositional data are vectors of non-negative components summing to a constant, typically equal 1, for simplicity. Their sample space is the standard simplex
$\mathbb{S}^{d}=\left\{(y_{1},...,y_{D})\ \big|\ y_{i}\geq 0,\sum_{i=1}^{D}y_{i}=1\right\},$ (1)
where $D$ denotes the number of variables (components), and d=D1d=D-1.
Examples of compositional data may be found in many different fields of study and the extensive scientific literature that has been published on the proper analysis of this type of data is indicative of its prevalence in real-life applications [^1] For a substantial number of specific examples of applications involving compositional data see (tsagris2020)..
It is unsurprising, considering the ubiquity of compositional data, that many applications of compositional data analysis incorporate covariates. Examples include glacial compositional data, household consumption expenditures, concentrations of chemical elements in soil samples, morphometric fish measurements, as well as data on elections, pollution, and energy, all of which are associated with some covariates. Beyond these cases, the literature provides numerous further applications of compositional regression. For example, oceanography research involving Foraminiferal compositions at various sea depths was analyzed in ait2003. In hydrochemistry, regression methods were used by otero2005 to distinguish anthropogenic from geological sources of river pollution in Spain. Economic studies such as morais2018 connected market shares to covariates, while political science research linked candidate vote percentages to relevant predictors (katz1999). In bioinformatics, compositional approaches have also been applied to microbiome data analysis (xia2013, chen2016, shi2016).
The practical demand for robust regression models tailored to compositional data has led to numerous methodological advances, especially in recent years. The first such model was introduced by ait2003—commonly known as Aitchison’s model—based on log-ratio transformations, yielding the log-ratio approach (LRA). ilr2003 advanced Aitchison’s model by applying an isometric log-ratio transformation. The stay-in-the-simplex approach on the other hand employs distributions and models defined on the simplex. Dirichlet regression for instance has been employed in compositional contexts gueorguieva2008, hijazi2009, melo2009. Moreover, iyengar2002 examined the generalized Liouville distribution family, which allows negative or mixed correlations and extends beyond Dirichlet distributions to include non-positive correlation structures. A less common approach is to ignore the compositional constraint and treat the data as though they were Euclidean, an approach termed raw data analysis (RDA) (baxter2001, baxter2005). A fourth approach is to employ a general family of transformations, namely the $\alpha$–transformation (tsagris2011) that interpolates between the and the RDA and the LRA, offers a higher flexibility and treats zero values naturally.
A limitation of the regression models discussed above is their inability to directly accommodate zero values. As a result, several models have been developed more recently to resolve this limitation. For instance, scealy2011 mapped compositional data onto the unit hyper-sphere and proposed the Kent regression, which naturally accounts for zeros. From a Bayesian perspective, spatial compositional data containing zeros were modeled in leininger2013. In the context of economics, mullahy2015 estimated regression models for share data where the proportions could assume zero values with non-negligible probability. Further econometric approaches suitable for handling zeros are reviewed in murteira2016. In addition, tsagris2015a introduced a regression framework based on minimizing the Jensen–Shannon divergence. tsagris2018 extended Dirichlet regression to allow zeros, resulting in what is termed zero-adjusted Dirichlet regression. More recently, alenazi2022 studied and examined the properties of the $\phi$-divergence regression models, which are suitable for compositional data with zeros.
When it comes spatial autocorrelation models, a simple version is the spatial distributed lag model with spatial lags on covariates, commonly known as the spatially lagged X (SLX) model. Unlike the general spatial Durbin or spatial autoregressive models, the SLX model incorporates spatial dependence only through the covariates, excluding the spatial lag of the dependent variable (lesage2009introduction, elhorst2014spatial).
Another model is the spatial autoregressive (SAR) model that similarly to the autoregressive model in time series it assumed that observations are influenced by those in nearby locations. In the SAR model, the dependent variable at each location is expressed as a function of both explanatory variables and a weighted average of the dependent variable values from neighboring areas. nguyen2021 proposed a SAR model that employs the ilr transformation.
A local form of linear regression, used to model spatially varying relationships, is the geographically weighted regression (GWR) is. Unlike traditional regression which assumes stationarity in the relationship between dependent and independent variables, GWR allows model parameters to vary over space. The integration of GWR with compositional data analysis is relatively recent. One key challenge is reconciling the spatial non-stationarity modeled by GWR with the constraints inherent in compositional data. Several approaches have been proposed. leininger2013 combined GWR with hierarchical Bayesian frameworks for compositional data with zero values, allowing for spatial priors that account for local variation. yoshida2021 applied the isometric log-ratio (ilr) transformation before applying GWR. This preserves the relative information between parts while enabling spatially varying coefficient estimation. Finally, clarotto2022 introduced a new power transformation, similar in spirit to the $\alpha$–transformation, for geostatistical modeling of compositional data.
[…]
2 The $\alpha$–regression
First the $\alpha$–transformation, used for the $\alpha$–regression, is defined, followed by the regression formulation.
2.1 The $\alpha$–transformation
tsagris2011 introduced the $\alpha$–transformation, a power-based mapping designed for compositional data, $\mathbf{y}=(y_{1},y_{2},\ldots,y_{D})$. For a given parameter $\alpha\in[-1,1]$, the transformation is defined in two steps. Each component is raised to the power $\alpha$ and renormalized to remain in the simplex
$\mathbf{u}=\left(\frac{y_{i}^{\alpha}}{\sum_{j=1}^{D}y_{j}^{\alpha}},\ldots,\frac{y_{D}^{\alpha}}{\sum_{j=1}^{D}y_{j}^{\alpha}}\right).$ (2)
This ensures $u=(u_{1},\ldots,u_{D})$ is itself a composition. To map compositions into Euclidean space for analysis, apply a linear transformation using the $D\times(D-1)$ Helmert sub-matrix $\mathbf{H}$:
$\mathbf{y}_{\alpha}=\frac{1}{\alpha}\left(D\mathbf{u}-1\right)\mathbf{H}^{\top},$ (3)
where $1$ denotes the $D$-dimensional vector of ones.
The transformation in Equation (3) is a one-to-one transformation which maps data inside the simplex onto a subset of $\mathbb{R}^{d}$ and vice versa for $\alpha\neq 0$. The corresponding sample space of Equation (3) is
$\mathbb{A}_{\alpha}^{d}=\left\{\mathbf{H}\mathbf{w}_{\alpha}(\mathbf{y})\ \big|-\frac{1}{\alpha}\leq w_{i,\alpha}\leq\frac{d}{\alpha},\sum_{i=1}^{d}w_{i,\alpha}=0\right\}.$ (4)
In effect, $y_{\alpha}$ which resembles a Box–Cox style mapping. The result $y_{\alpha}$ is an unconstrained vector in Euclidean space, suitable for standard multivariate statistical techniques. When $\alpha=1$, the transformation corresponds (up to scaling) to raw data analysis (RDA). When $\alpha=-1$, the transformation is aligned with RDA as well, but using the inverse of the compositional data. As $\alpha\to 0$, the transformation converges to the ilr transformation used in log-ratio analysis (LRA)
$\mathbf{y}_{0}=\left(\log\left({\frac{y_{1}}{\prod_{j=1}^{D}x_{j}^{1/D}}}\right),\ldots,\log\left({\frac{y_{D}}{\prod_{j=1}^{D}y_{j}^{1/D}}}\right)\right)\mathbf{H}^{\top}.$ (5)
Thus, the $\alpha$–transformation provides a continuum between RDA and LRA, allowing analysts to choose the most appropriate representation of compositional data based on empirical performance or theoretical considerations.
2.2 The $\alpha$–regression
The $\alpha$–regression has the potential to improve the regression predictions with compositional data by adapting the $\alpha$–transformation to the data set’s geometry. We assume that the conditional mean of the observed composition can be written as a non-linear function of some covariates
$\mu_{i}=\begin{cases}\displaystyle\frac{1}{1+\sum_{j=1}^{D}e^{\mathbf{x}^{\top}\mathbf{\beta}_{j}}}&\text{for }i=1\\[12.91663pt] \displaystyle\frac{e^{\mathbf{x}^{\top}\mathbf{\beta}_{i}}}{1+\sum_{j=1}^{D}e^{\mathbf{x}^{\top}\mathbf{\beta}_{j}}}&\text{for }i=2,\ldots,D\end{cases}$ (6)
where
$\boldsymbol{\beta}_{i}=\left(\beta_{0i},\beta_{1i},...,\beta_{pi}\right)^{\top},\ i=1,...,d\ \ \text{and $p$ denotes the number of covariates}.$
tsagris2015b used the log-likelihood of the multivariate normal distribution, but in this paper the regression is formulated as a non-linear least squares problem, where the minimizing function is the sum of squares of the errors (SSE)
$\text{SSE}\left(\mathbf{Y},\mathbf{X};\alpha,\mathbf{B}\right)=\sum_{i=1}^{n}\|\mathbf{y}_{i,\alpha}-\mathbf{\mu}_{i,\alpha}\|_{2}=\sum_{i=1}^{n}\left(\mathbf{y}_{i,\alpha}-\mathbf{\mu}_{i,\alpha}\right)^{\top}\left(\mathbf{y}_{i,\alpha}-\mathbf{\mu}_{i,\alpha}\right),$ (7)
where $\mathbf{y}_{i,\alpha}$ and $\mathbf{m}_{i,\alpha}$ are the $\alpha$–transformations applied to the $i$-th response and fitted compositional vectors, respectively and $\|.\|_{2}$ denotes the $L_{2}$ norm. Application of the stay-in-the-simplex power transformation (2) to the fitted vectors yields a simplified expression
$\frac{\mu_{i}^{\alpha}}{\sum_{j=1}^{D}\mu_{j}^{\alpha}}=\frac{\left(\frac{e^{{\bf x}^{\top}\boldsymbol{\beta}_{i}}}{1+\sum_{j=1}^{D}e^{{\bf x}^{\top}\boldsymbol{\beta}_{j}}}\right)^{\alpha}}{\frac{1+\sum_{k=1}^{D}\left(e^{{\bf x}^{\top}\boldsymbol{\beta}_{k}}\right)^{\alpha}}{\left(1+\sum_{j=1}^{D}e^{{\bf x}^{\top}\boldsymbol{\beta}_{j}}\right)^{\alpha}}}=\frac{\left(e^{{\bf x}^{\top}\boldsymbol{\beta}_{i}}\right)^{\alpha}}{1+\sum_{j=1}^{D}\left(e^{{\bf x}^{\top}\boldsymbol{\beta}_{i}}\right)^{\alpha}}.$
For a given value of $\alpha$, the matrix of the regression coefficients $\mathbf{B}=\left(\mathbf{\beta}_{1},\ldots,\mathbf{\beta}_{d}\right)$ is estimated using a modification of the Levenberg-Marquardt algorithm [^2] This algorithm interpolates between the Gauss–Newton algorithm and the method of gradient descent.. The R package minpack.lm (minpacklm2023) is employed to this end [^3] The relevant gradient vector, and the Hessian matrix are provided in the Appendix. The Newton-Raphson algorithm was also tested but it is slower..
2.2.1 Limiting case of $\alpha\rightarrow 0$
tsagris2016 presented the proof that as $\alpha\rightarrow 0$, the $\alpha$–transformation (3) converges to the ilr transformation (5). Following similar calculations one can show that
$\lim_{\alpha\rightarrow 0}{\frac{1}{\alpha}\left(D\frac{\mu_{i}^{\alpha}}{\sum_{j=1}^{D}\mu_{j}^{\alpha}}-1\right)}\rightarrow\mathbf{x}\beta_{i}-\frac{\sum_{j=1}^{D}\mathbf{x}\beta_{j}}{D},$
which corresponds to the regression after the centered log-ratio transformation [the ilr transformation (5) without the right multiplication by the Helmert matrix]. This implies that there are $D$ vectors of $\mathbf{\beta}$ regression coefficients. But, since the first set of regression coefficients equals zero, if we subtract this vector from the rest of the $\mathbf{\beta}$ vectors we end up with the regression coefficients of the additive log-ratio (alr) regression
$\log\left(\frac{y_{i}}{y_{1}}\right)=\mathbf{x}^{\top}\boldsymbol{\beta}_{i},\ \ i=2,\ldots,D$
2.2.2 Choosing $\alpha$
In the regression setting the optimal value of $\alpha$ is data-driven, and there are two ways to estimate its value. The first is minimize the Kullback-Leibler divergence (KLD) between the observed and fitted compositions $\text{KLD}\left(\mathbf{y},\mathbf{\mu}\right)=\sum_{i=1}^{n}\sum_{j=1}^{D}\mathbf{y}_{ij}\log{y_{i}j/\mu_{ij}}$. This results in a double minimization problem. For a given value of $\alpha$ one must minimize the SSE (7) in order to obtain the regression coefficients and then minimize the KLD with respect to $\alpha$ to obtain the optimal value of $\alpha$. This way and with the choice of the KLD, the value of $\alpha$ is independent of the SSE, since the SSE is not comparable across the different values of $\alpha$.
The second option is to examine $\alpha$ as a hyper-parameter whose value is chosen by minimizing the KLD via CV, e.g. 10-fold CV. (tsagris2015b).
2.2.3 Marginal effects
To account for the difficult interpretation of the regression coefficients, the marginal effects (MEs), given below, may be used
$\text{ME}_{k}=\frac{\partial\mu_{i}}{\partial x_{k}}=\left\{\begin{array}[]{ll}-\mu_{1}\sum_{j=1}^{d}\beta_{jk}\mu_{j+1}&\text{for}\ \ i=1\\ \mu_{i}\left(\beta_{i-1,k}-\sum_{j=1}^{d}\beta_{jk}\mu_{j+1}\right)&\text{for}\ \ i=2,\ldots,D\end{array}\right\},$ (10)
where $\sum_{i=1}^{D}\frac{\partial\mu_{i}}{\partial x_{k}}=0$, because $\sum_{i=1}^{D}\mu_{i}=1$. The sum of the MEs sums to zero, because if all components increase, one at least component must decrease by the same amount so that the unity sum constraint is preserved.
The average MEs (AME) across all observations are then computed as
$\text{AME}_{k}=\frac{1}{n}\sum_{i=1}^{n}\frac{\partial\mu_{i}}{\partial x_{k}}.$
Standard errors can be computed via bootstrap or the delta method, accounting for estimation uncertainty in both $\widehat{\mathbf{\beta}}$, $\widehat{\mathbf{\gamma}}$, and $\widehat{\mathbf{\mu}}$.
2.2.4 Standard error of the MEs
The covariance matrix of MEs is derived using the delta method. Let $\mathbf{\theta}=\text{vec}(\mathbf{B})$ be the $dp\times 1$ vector of stacked regression coefficients, where $\mathbf{B}=(\mathbf{\beta}_{1},\ldots,\mathbf{\beta}_{d})$ is the $d\times p$ matrix of coefficients.
The general method of the Delta method reads that for a function $g(\mathbf{\theta})$ of the parameters:
$\text{Var}(g(\mathbf{\theta}))\approx\mathbf{J}\text{Var}(\mathbf{\theta})\mathbf{J}^{\top}$
where $\mathbf{J}=\frac{\partial g}{\partial\mathbf{\theta}^{\top}}$ is the Jacobian matrix.
For observation $i$, component $\ell$, and covariate $k$, the ME is:
$m_{i\ell k}=\frac{\partial\mu_{i\ell}}{\partial x_{ik}}$ (11)
First, we need the derivatives of $\mu$ with respect to $\mathbf{\beta}$:
$\frac{\partial\mu_{ij}}{\partial\beta_{rs}}=\begin{cases}-\mu_{i1}\mu_{ir+1}x_{is}&\text{if }j=1\\[4.30554pt] \mu_{ij}(1-\mu_{ij})x_{is}&\text{if }j=r+1\\[4.30554pt] -\mu_{ij}\mu_{ir+1}x_{is}&\text{otherwise},\end{cases}$
where $r\in\{1,\ldots,d\}$ and $s\in\{1,\ldots,p\}$.
The Jacobian of the ME for the reference component ($\ell=1$) is
$\frac{\partial}{\partial\beta_{ms}}\left(\frac{\partial\mu_{i\ell}}{\partial x_{ik}}\right)=\begin{cases}-\delta_{sk}\mu_{i1}\mu_{im+1}-\frac{\partial\mu_{i1}}{\partial\beta_{ms}}\sum_{j=1}^{d}\beta_{jk}\mu_{ij+1}-\mu_{i1}\sum_{j=1}^{d}\beta_{jk}\frac{\partial\mu_{ij+1}}{\partial\beta_{ms}}&\text{ for }\ell=1\\ \delta_{sk}\delta_{m,\ell-1}\mu_{i\ell}-\delta_{sk}\mu_{i\ell}\mu_{im+1}+\frac{\partial\mu_{i\ell}}{\partial\beta_{ms}}\left[\beta_{\ell-1,k}-\sum_{j=1}^{d}\beta_{jk}\mu_{ij+1}\right]-\\ \mu_{i\ell}\sum_{j=1}^{d}\beta_{jk}\frac{\partial\mu_{ij+1}}{\partial\beta_{ms}}&\text{ for }\ell=2,\ldots,D,\end{cases}$
where $\delta_{sk}$ is the Kronecker delta ($\delta_{sk}=1$ if $s=k$, 0 otherwise).
For observation $i$, component $\ell$, and covariate $k$, the variance is
$\text{Var}\left(\frac{\partial\mu_{i\ell}}{\partial x_{ik}}\right)=\mathbf{J}_{i\ell k}\,\text{Var}(\mathbf{\theta})\,\mathbf{J}_{i\ell k}^{\top},$ (12)
and the covariance matrix of the AMEs is
$\text{Cov}(\text{AME})=\left(\frac{1}{n}\sum_{i=1}^{n}\mathbf{J}_{i}\right)\text{Var}(\mathbf{\theta})\left(\frac{1}{n}\sum_{i=1}^{n}\mathbf{J}_{i}\right)^{\top}.$ (13)
2.2.5 Advantages and limitations of the $\alpha$–regression
The advantages of the $\alpha$–regression are: a) ability to handle zeros naturally without imputation. b) Flexible, as $\alpha$ provides a continuum from power transforms to log-ratio methods. c) Often yields better predictive performance than classical methods. d) This method balances the strengths of power transformations and log-ratio methods, providing a flexible and effective tool for predictive modeling on the simplex. Disadvantages on the other hand are a) the interpretability of regression coefficients is reduced compared to log-ratio approaches. b) The focus is mainly on prediction rather than inference; theoretical properties of estimators have not been rigorously developed.
2.3 Extension of the $\alpha$–regression and related models
. The $\alpha$–regression is based upon minimization of the $L_{2}$ norm (7). In a similar fashion we can minimize the $L_{1}$ norm, yielding the $\alpha$–minium absolute deviations ($\alpha$–MAD) regression
$\text{MAD}\left(\mathbf{Y},\mathbf{X};\alpha,\mathbf{B}\right)=\sum_{i=1}^{n}\|\mathbf{y}_{i,\alpha}-\mathbf{\mu}_{i,\alpha}\|_{1}=\sum_{i=1}^{n}\sum_{j=1}^{d}|\mathbf{y}_{ij,\alpha}-\mathbf{\mu}_{ij,\alpha}|.$ (14)
Following tsagris2025 we formulated the $\alpha$–MAD regression as a univariate regression problem, by using the vectorization operation for the responses and by constructing the design matrix in a suitable manner. To make the estimation efficient we utilized the command nlrq() from the R package quantreg (quantreg2025). This approach exhibits dependence on initialization and does not ensure convergence while use of the R built in functions optim or nlminb is computationally more expensive. For these reasons we decided to not pursue this approach further.
[…]
3.1 The $\alpha$–SLX model
The $\alpha$–SLX model extends the standard $\alpha$-regression by incorporating spatial spillover effects through the covariates. The fitted compositional values are given by:
$\mu_{i}=\begin{cases}\displaystyle\frac{1}{1+\sum_{j=1}^{D}e^{\mathbf{x}^{\top}\mathbf{\beta}_{j}+(\mathbf{W}\mathbf{x})^{\top}\mathbf{\gamma}_{j}}}&\text{for }i=1\\[12.91663pt] \displaystyle\frac{e^{\mathbf{x}^{\top}\mathbf{\beta}_{i}+(\mathbf{W}\mathbf{x})^{\top}\mathbf{\gamma}_{i}}}{1+\sum_{j=1}^{D}e^{\mathbf{x}^{\top}\mathbf{\beta}_{j}+(\mathbf{W}\mathbf{x})^{\top}\mathbf{\gamma}_{j}}}&\text{for }i=2,\ldots,D.\end{cases}$ (15)
The matrices of regression coefficients $\mathbf{B}=\left(\mathbf{\beta}_{1},\ldots,\mathbf{\beta}_{d}\right)$ and $\mathbf{\Gamma}=\left(\mathbf{\gamma}_{1},\ldots,\mathbf{\gamma}_{d}\right)$ in the same way as in the $\alpha$–regression.
3.1.1 The contiguity matrix
The Euclidean distance between any two pairs of latitude and longitude, $(\nu_{i},v_{i})$ and $(\nu_{j},v_{j})$. As mentioned earlier, the locations are first mapped from their polar to their Cartesian coordinates (after transforming the degrees into radians)
$\mathbf{c}_{i}=\left(\cos(\nu_{i}),sin(\nu_{i})\cos(v_{i}),\sin(\nu_{i})\sin(v_{i})\right)\ \text{and}\ \mathbf{c}_{j}=\left(\cos(\nu_{j}),sin(\nu_{j})\cos(v_{j}),\sin(\nu_{j})\sin(v_{j})\right).$
The Euclidean distance between $\mathbf{c}_{i}$ and $\mathbf{c}_{j}$ is
$d(\mathbf{c}_{i},\mathbf{c}_{j})=d_{ij}^{2}=\left\|\mathbf{c}_{i}-\mathbf{c}_{j}\right\|^{2}=\left\|\mathbf{c}_{i}\right\|^{2}+\left\|\mathbf{c}_{j}\right\|^{2}-2\mathbf{c}_{i}^{\top}\mathbf{c}_{j}=2\left(1-\mathbf{c}_{i}^{\top}\mathbf{c}_{j}\right).$
For the $i$-th location, compute the region with the the $k$ nearest neighbors $\mathcal{C}_{ik}$ and zero the rest, that is
$\tilde{w}_{ij}=\begin{cases}1/d_{ij}^{2}&\text{if}\ j\in\mathcal{C}_{ik}\\ \tilde{w}_{ij}=0&\text{else}.\end{cases}$ (16)
The $(i,j)$ elemets of the contiguity matrix $\mathbf{W}$ are then defined as $w_{ij}=\tilde{w}_{ij}/\sum_{j=1}^{n}\tilde{w}_{ij}$.
3.1.2 Choosing $\alpha$ and $k$
The choice of the optimal values of $\alpha$ and of $k$ is again data-driven and can be performed via CV, but this time the spatial 10-fold cross-validation CV protocol is employed, where the metric of performance is again the KLD.
The R package blockCV (blockcv2019) implements spatial cross-validation techniques designed to address the spatial autocorrelation inherent in geographical data. Unlike the traditional 10-fold CV, which can lead to overly optimistic model performance estimates when data points are spatially clustered (roberts2017), the spatial version partitions data into spatially separated training and testing folds. This ensures that the testing data are spatially independent from the training data, providing more realistic assessments of model generalization to new geographic areas.
3.1.3 Spatial MEs
The spatial MEs (SMEs) consist of three components, the direct, the indirect and the total. The direct MEs measure the impact of a change in the local covariate $x_{k}$ on the local composition component $\mu_{i}$. The following formulas are identical to the standard $\alpha$–regression MEs (10), as they depend only on the $\mathbf{\beta}$ coefficients and do not involve spatial terms.
The direct SMEs measure the change in the covariate values
$DE=\frac{\partial\mu_{i}}{\partial x_{k}}=\begin{cases}-\mu_{1}\sum_{j=1}^{d}\beta_{jk}\mu_{j+1}&\text{for }i=1\\ \mu_{i}\left(\beta_{i-1,k}-\sum_{j=1}^{d}\beta_{jk}\mu_{j+1}\right)&\text{for }i=2,\ldots,D.\end{cases}$ (17)
The indirect (spillover) SMEs measure the impact of a change in the spatially lagged covariate $(\mathbf{W}\mathbf{x})_{k}$ (i.e., the weighted average of neighboring values) on the local composition component $\mu_{i}$. They have the same functional form as the direct effects, with $\mathbf{\gamma}$ replacing $\mathbf{\beta}$. This structural symmetry reflects how spatial spillovers operate through the same multiplicative mechanism as direct effects.
$IDE=\frac{\partial\mu_{i}}{\partial(\mathbf{W}\mathbf{x})_{k}}=\begin{cases}-\mu_{1}\sum_{j=1}^{d}\gamma_{jk}\mu_{j+1}&\text{for }i=1\\ \mu_{i}\left(\gamma_{i-1,k}-\sum_{j=1}^{d}\gamma_{jk}\mu_{j+1}\right)&\text{for }i=2,\ldots,D.\end{cases}$ (18)
The total SMEs combine both direct and indirect SMEs representing the full impact of a simultaneous change in both local and neighboring covariate values.
$\frac{\partial\mu_{i}}{\partial x_{k}}+\frac{\partial\mu_{1}}{\partial(\mathbf{W}\mathbf{x})_{k}}=\begin{cases}-\mu_{1}\sum_{j=1}^{d}(\beta_{jk}+\gamma_{jk})\mu_{j+1}&\text{for }i=1\\ \mu_{i}\left[(\beta_{i-1,k}+\gamma_{i-1,k})-\sum_{j=1}^{d}(\beta_{jk}+\gamma_{jk})\mu_{j+1}\right]&\text{for }i=2,\ldots,D.\end{cases}$ (19)
3.1.4 Properties of the SMEs
Some properties regarding the SMEs are delineated below.
  • The sum of the SMEs across all components equals zero:
    $\sum_{i=1}^{D}\frac{\partial\mu_{i}}{\partial x_{k}}=0\quad\text{and}\quad\sum_{i=1}^{D}\frac{\partial\mu_{i}}{\partial(\mathbf{W}\mathbf{x})_{k}}=0$ (20)
    This ensures that the composition remains on the simplex after perturbations.
  • All SMEs depend on the current composition values $\mathbf{\mu}$, making them observation-specific and state-dependent.
  • Direct and indirect effects share the same functional form, differing only in the coefficient vectors used ($\mathbf{\beta}$ vs $\mathbf{\gamma}$).
  • The spatial weights matrix $\mathbf{W}$ determines which neighbors contribute to spillover effects. We remind that row-standardization is used such that $\sum_{j}w_{ij}=1$.
  • The standard error of the SMEs can be computed in a manner similar to the MEs of the $\alpha$–regression and the formulas of their Jacobians can be found in Appendix D.
3.1.5 Prediction of new values
To predict the compositions for new observations $\mathbf{x}_{new}$, we must first construct the matrix $\mathbf{W}_{new}$ which contains the row normalized distances from the new locations to the existing ones, and then use the following formula
$\widehat{\mu}_{i}=\begin{cases}\displaystyle\frac{1}{1+\sum_{j=1}^{D}e^{\mathbf{x}_{new}^{\top}\mathbf{\beta}_{j}+(\mathbf{W}_{new}\mathbf{x})^{\top}\mathbf{\gamma}_{j}}}&\text{for }i=1\\[12.91663pt] \displaystyle\frac{e^{\mathbf{x}_{new}^{\top}\mathbf{\beta}_{i}+(\mathbf{W}_{new}\mathbf{x})^{\top}\mathbf{\gamma}_{i}}}{1+\sum_{j=1}^{D}e^{\mathbf{x}_{new}^{\top}\mathbf{\beta}_{j}+(\mathbf{W}_{new}\mathbf{x})^{\top}\mathbf{\gamma}_{j}}}&\text{for }i=2,\ldots,D.\end{cases}$ (21)
3.2 The $\alpha$–SAR model
Inspired by the SAR for multinomial regression we define the following formulation
$p_{i}$ $=$ $\rho\mathbf{W}p_{i}+\mathbf{x}^{\top}\mathbf{\beta}_{i}+\epsilon_{i}$
$(\mathbf{I}_{n}-\rho\mathbf{W})\,p_{i}$ $=$ $\mathbf{x}^{\top}\mathbf{\beta}_{i}+\epsilon_{i}$
$p_{i}$ $=$ $S\left(\rho\right)^{-1}\mathbf{x}^{\top}\mathbf{\beta}_{i}+S\left(\rho\right)^{-1}\epsilon_{i},$
for $i=1,\dots,D$, where $\rho\in(-1,1)$ is the spatial autoregressive parameter measuring spillover strength, $S\left(\rho\right)=\mathbf{In}-\rho\mathbf{W}$ is the spatial multiplier matrix, and $\epsilon_{i}\sim\mathcal{N}(\mathbf{0},\sigma_{i}^{2}\mathbf{I}_{n})$.
The fitted values are then defined in the same manner as in (6)
$\mu_{i}=\begin{cases}\displaystyle\frac{1}{1+\sum_{j=1}^{D}e^{(\mathbf{I}_{n}-\rho\mathbf{W})^{-1}\mathbf{x}^{\top}\mathbf{\beta}_{j}}}&\text{for }i=1\\[12.91663pt] \displaystyle\frac{e^{(\mathbf{I}_{n}-\rho\mathbf{W})^{-1}\mathbf{x}^{\top}\mathbf{\beta}_{i}}}{1+\sum_{j=1}^{D}e^{(\mathbf{I}_{n}-\rho\mathbf{W})^{-1}\mathbf{x}^{\top}\mathbf{\beta}_{j}}}&\text{for }i=2,\ldots,D.\end{cases}$ (22)
[…]
3.2.1 Computational challenges
[…]
3.2.2 SMEs
The direct effects measure the impact of a change in location $i$’s covariate on location $i$’s own composition:
$\frac{\partial\mu_{i\ell}}{\partial x_{ik}}=\begin{cases}-\mu_{i1}\sum_{j=1}^{d}\beta_{jk}\mu_{ij+1}\cdot[S(\rho)^{-1}]_{ii}&\text{for }\ell=1\\[8.61108pt] \mu_{i\ell}\left[\beta_{\ell-1,k}-\sum_{j=1}^{d}\beta_{jk}\mu_{ij+1}\right]\cdot[S(\rho)^{-1}]_{ii}&\text{for }\ell=2,\ldots,D.\end{cases}$ (23)
The indirect effect at location $i$ (summing spillovers from all neighbors) is
$\frac{\partial\mu_{i\ell}}{\partial x_{jk}}=\begin{cases}-\mu_{i1}\sum_{j=1}^{d}\beta_{jk}\mu_{ij+1}\cdot\sum_{j\neq i}[S(\rho)^{-1}]_{ij}&\text{for }\ell=1\\[8.61108pt] \mu_{i\ell}\left[\beta_{\ell-1,k}-\sum_{j=1}^{d}\beta_{jk}\mu_{ij+1}\right]\cdot\sum_{j\neq i}[S(\rho)^{-1}]_{ij}&\text{for }\ell=2,\ldots,D\end{cases}$
The total SMEs are the sum of the direct and indirect effects.
3.2.3 Prediction of new values
Denote the new $m$ covariate values by $\mathbf{X}^{new}$ located at new, unseen in the model, coordinates. We stack the the new covariate values under the observed ones to create the augmented design matrix
$\mathbf{X}^{aug}=\left(\begin{array}[]{cc}\mathbf{X}&\mathbf{X}^{new}\end{array}\right).$
Similarly define
$\mathbf{W}^{aug}=\left(\begin{array}[]{cc}\mathbf{W}&\mathbf{W}^{new}\\ \mathbf{W}^{new}&\mathbf{W}\end{array}\right)$
to be the augmented contiguity matrix, where $\mathbf{W}^{new}$ denotes the distances of the new locations from the observed ones. Note that $\mathbf{W}^{aug}$ is row standardised. $\mathbf{X}^{aug}$ contains $n+m$ rows, and $\mathbf{W}^{aug}$ is of dimensions $(n+m)\times(n+m)$, where $n$ is the sample size of the observed sample, upon which the estimates are derived.
The predicted values are given by
$\widehat{y}^{aug}_{ij}=\begin{cases}\displaystyle\frac{1}{1+\sum_{\ell=1}^{D}e^{(\mathbf{I}_{n+m}-\rho\mathbf{W}^{aug})^{-1}\left(\mathbf{x}_{j}^{aug}\right)^{\top}\mathbf{\beta}_{\ell}}}&\text{for }i=1\\[12.91663pt] \displaystyle\frac{e^{(\mathbf{I}_{n+m}-\rho\mathbf{W}^{aug})^{-1}\left(\mathbf{x}_{j}^{aug}\right)^{\top}\mathbf{\beta}_{i}}}{1+\sum_{\ell=1}^{D}e^{(\mathbf{I}_{n+m}-\rho\mathbf{W}^{aug})^{-1}\left(\mathbf{x}_{j}^{aug}\right)^{\top}\mathbf{\beta}_{\ell}}}&\text{for }i=2,\ldots,D.\end{cases}$
Stacking the predicted values, in a matrix format, $\widehat{\mathbf{Y}}^{aug}=\left(\begin{array}[]{c}\widehat{\mathbf{Y}}\\ \widehat{\mathbf{Y}}^{new}\end{array}\right),$ we observe that we the predictions for the new covariate values at the new locations are placed in the bottom $m$ rows of $\widehat{\mathbf{Y}}^{aug}$.
3.2.4 A non-linear spatial autocorrelation test
Moran’s I test (moran1950) is the oldest and most widespread test of spatial autocorrelation. In the univariate case, the test statistic is basically the Pearson correlation between the response $\mathbf{y}$ and the spatially weighted average $\mathbf{Wy}$. The p-value for the hypothesis of no spatial correlation can be computed asymptotically or via permutations.
[…]
$SSE\left(\mathbf{Y},\mathbf{X};\alpha,h,\mathbf{B}\right)=\sum_{i=1}^{n}\left(\mathbf{y}_{i,\alpha}-\mathbf{\mu}_{i,\alpha}\right)^{\top}\mathbf{W}_{i}\left(\mathbf{y}_{i,\alpha}-\mathbf{\mu}_{i,\alpha}\right),$ (26)
where $\mathbf{W}_{i}=\text{diag}\left\{w_{i1},\ldots,w_{in}\right\}$, is the weighting matrix corresponding to the weights allocated to each observation. A common weighting function is the Gaussian kernel
$w_{ij}=\exp\left(-\frac{d_{ij}^{2}}{2h^{2}}\right),$ (27)
where $d_{ij}$ is the distance between location $i$ and $j$, and $h$ is the bandwidth parameter controlling the degree of spatial smoothing.
[…]
3.3.1 Computing $d_{ij}$ in the weighting scheme
Some researchers tend to compute the Euclidean distance between two pairs of latitude and longitude, $(\nu_{i},v_{i})$ and $(\nu_{j},v_{j})$, $d_{ij}=\sqrt{\left(\nu_{i}-\nu_{j}\right)^{2}+\left(v_{i}-v_{j}\right)^{2}}$. There is a fundamental flaw with this approach which is highlighted by mardia2000, pg. 13. Consider for instance the case of two coordinates whose latitude (or longitude) values are $359^{\circ}$ and $1^{\circ}$. Using the previous naive approach yields a distance between the two values $359^{\circ}-1^{\circ}=358^{\circ}$, but the actual distance between them is only $2^{\circ}$. To account for this, the pair of coordinates must first be transformed into their Euclidean coordinates, prior to the application of the Euclidean distance.
3.3.2 Choice of $\alpha$ and $h$
[…]
3.3.3 Computational details
  • Similarly to the $\alpha$–regression, the stay-in-the-simplex power transformation (2) is written as
    $\frac{\mu_{i}^{\alpha}}{\sum_{j=1}^{D}\mu_{j}^{\alpha}}=\frac{\left(e^{\mathbf{x}^{\top}\boldsymbol{\beta}_{i}}\right)^{\alpha}}{1+\sum_{j=1}^{D}\left(e^{\mathbf{x}^{\top}\boldsymbol{\beta}_{i}}\right)^{\alpha}}=\frac{\left(e^{\alpha\mathbf{x}^{\top}\boldsymbol{\beta}_{i}}\right)}{1+\sum_{j=1}^{D}\left(e^{\alpha\mathbf{x}^{\top}\boldsymbol{\beta}_{i}}\right)}.$
  • The the weighting function (27) becomes $w_{ij}=\exp\left(-\frac{d_{ij}^{2}}{2h^{2}}\right)=\exp\left(\frac{\mathbf{c}_{i}^{\top}\mathbf{c}_{j}-1}{h^{2}}\right)$.
  • The minimization of the $SSE$ takes place for specific values of $\alpha$ and $h$. When passing the arguments of the $SSE$ in the command minpack.lm::nls.lm(), the quantity $\alpha\mathbf{x}$ is pre-computed and passed as an argument.
  • The function minpack.lm::nls.lm() requires a function that outputs the residuals. So, in order to perform weighted lest squares we multiply the weights $\mathbf{w}_{i}$ by the residuals $\mathbf{r}_{i}$.
  • For each observation $i$, we can compute the regression coefficients for different values of $h$. This is useful during the cross-validation protocol.
3.3.4 SMEs
[…]
$\frac{\partial\mu_{1}\left(\nu_{i},v_{i}\right)}{\partial x_{k}}=\begin{cases}-\mu_{1}\left(\nu_{i},v_{i}\right)\sum_{j=1}^{d}\beta_{jk}\left(\nu_{i},v_{i}\right)\mu_{j+1}\left(\nu_{i},v_{i}\right)&\text{for }i=1\\ \mu_{\ell}\left(\nu_{i},v_{i}\right)\left[\beta_{i-1,k}\left(\nu_{i},v_{i}\right)-\sum_{j=1}^{d}\beta_{jk}\left(\nu_{i},v_{i}\right)\mu_{j+1}\left(\nu_{i},v_{i}\right)\right]&\text{for }\ell=2,\ldots,D.\end{cases}$ (28)
Just like in the $\alpha$–regression, the $\sum_{\ell=1}^{D}\frac{\partial\mu_{\ell}\left(\nu_{i},v_{i}\right)}{\partial x_{k}}=0$, but this time, this is true for every location.
Jacobian vector at location $(\nu_{i},u_{i})$
Define $\mathbf{\theta}_{i}=\text{vec}(\mathbf{B}_{i})\in\mathbb{R}^{dp}$, where $\mathbf{B}_{i}=(\mathbf{\beta}_{i,1},\ldots,\mathbf{\beta}_{i,d})$ is the $d\times p$ matrix the of location-specific coefficients (at $(\nu_{i},u_{i})$). For observation $i$, component $\ell$, and covariate $k$, at location $(\nu_{i},u_{i})$ the MEs are defined as
$\frac{\partial\mu_{i\ell k}}{\partial\beta_{i,ms}}=\begin{cases}-\delta_{sk}\mu_{i1}\mu_{im+1}-\frac{\partial\mu_{i1}}{\partial\beta_{i,ms}}\sum_{j=1}^{d}\beta_{i,jk}\mu_{ij+1}-\mu_{i1}\sum_{j=1}^{d}\beta_{i,jk}\frac{\partial\mu_{ij+1}}{\partial\beta_{i,ms}}&\text{for }\ell=1\\ \delta_{sk}\delta_{m,\ell-1}\mu_{i\ell}-\delta_{sk}\mu_{i\ell}\mu_{im+1}+\frac{\partial\mu_{i\ell}}{\partial\beta_{i,ms}}\left(\beta_{i,\ell-1,k}-\sum_{j=1}^{d}\beta_{i,jk}\mu_{ij+1}\right)\\ -\mu_{i\ell}\sum_{j=1}^{d}\beta_{i,jk}\frac{\partial\mu_{ij+1}}{\partial\beta_{i,ms}}&\text{for }\ell=2,\ldots,D,\end{cases}$
where the derivatives of the composition with respect to the location-specific coefficients are:
$\frac{\partial\mu_{ip}}{\partial\beta_{i,rs}}=\begin{cases}-\mu_{i1}\mu_{ir+1}x_{is}&\text{if }p=1\\[4.30554pt] \mu_{ip}(1-\mu_{ip})x_{is}&\text{if }p=r+1\\[4.30554pt] -\mu_{ip}\mu_{ir+1}x_{is}&\text{otherwise},\end{cases}$
where $r\in\{1,\ldots,d\}$ and $s\in\{1,\ldots,p\}$.
4 Application to real data sets
Real-data applications show that the $\alpha$–regression can outperform the standard log-ratio-based regression, in terms of predictive performance, particularly when zeros are present, which can be further improved by taking into account the spatial dependencies.
[…]
4.1 Agricultural economics data set
Data regarding crop productivity in the Greek NUTS II region of Thessaly during the 2017-218 cropping year were supplied by the Greek Ministry of Agriculture, also known as farm accountancy data network (FADN) data. The data refer to a sample of farms and initially they consisted of 20 crops, but after grouping and aggregation they were narrowed down to 5 crops [^6] A larger version of this data set was used in mattas2025. Following the EU Regulation No1166/2008 that establishes a framework for European statistics at the level of agricultural holdings the aggregation took place across different output of crops.. These crops are Cereals, Cotton, Tree crops, Other annual crops and pasture and Grapes and wine. For each of the 168 farms with unique coordinates, the cultivated area in each of these 5 grouped crops is known. Note that due to the existence of zeros, the LRA approach, i.e. the family of the $\alpha$–regression models presented earlier with $\alpha=0$ is not applicable.
Figure 1(a) shows the location of Thessaly region in Greece, and Figure 1(b) shows the locations of the farms. The majority of the farms cultivate cereals and only few farms hold grapes and wine. Specifically, 84.52% of the farms cultivate cereals, 50.00% cultivate Cotton, 40.48% maintain tree crops, 81.55% hold other annual crops and pasture, and finally only 16.67% of the farms own grapes and wine.
Refer to caption Refer to caption
(a) Region of Thessaly within Greece. (b) The locations of the 168 farms.
Figure 1: The Thessaly region in Greece.
The goal is to examine the relationship between the composition of the cultivated area and the The covariates were the following:
  • Human Influence Index (HII, direct human influence on ecosystems). Zero value represents no human influence and 64 represents maximum human influence possible. The index uses all 8 measurements of human presence: Population Density/km2, Score of Railroads, Score of Major Roads, Score of Navigable, Rivers, Score of Coastlines, Score of Nighttime Stable Lights Values, Urban Polygons, Land Cover Categories. The range of observed values is 16.0846.6916.08-46.69, with an average of 29.021.
  • The soil pH (CaCl2). The range of values observed was between 06.990-6.99 and the average was 6.33.
  • Topsoil organic carbon content (SOC). The content (%) in the surface horizon of soils. The values ranged from 0.54 up to 10.07 with an average equal to 1.41.
[…]
Table 1: Agricultural economics data set: average results regarding the optimal choice of $\alpha$, $k$, $h$, KLD and running time (in seconds) for each of the three models.
Model KLD $\alpha$ $k$ $h$ Running time
$\alpha$–regression 0.810 0.775 3.308
$\alpha$–SLX 1.603 0.550 6 172.860
$\alpha$–SAR 0.608 1.000 5 1258.929
[…]
We then fitted the regression models using the optimal parameters obtained based on the CV protocol and computed the correlations (component-wise) between the observed and fitted compositions. Table 2 contains these correlations. Evidently, the $\alpha$–SAR model seems to have achieved the best fit. It is important to highlight that the spatial autoregressive parameter $\rho$ of the $\alpha$–SAR model was equal to -0.148 with a standard error equal to 0.0681. A simple $t$-test shows that the coefficient is statistically significant.
Table 2: Agricultural economics data set: Pearson correlations between each pair of the observed and fitted components for each of the three regression models.
Cereals Cotton Tree crops Other annual crops Grapes and wine
and pasture
$\alpha$–regression 0.324 0.589 0.603 0.348 0.224
$\alpha$–SLX 0.327 0.626 0.627 0.413 0.328
$\alpha$–SAR 0.318 0.582 0.608 0.357 0.230
[…]
4.2 Meuse river data set
This data set gives locations and topsoil heavy metal concentrations, along with a number of soil and landscape variables at the observation locations, collected in a flood plain of the river Meuse, near the village of Stein (Netherlands). Heavy metal concentrations are from composite samples of a squared area of approximately 15m $\times$ 15m. There are measurements (all measured in mg kg-1 (ppm)): topsoil cadmium concentration (zero cadmium values in the original data set have been shifted to 0.2 (half the lowest non-zero value)), topsoil copper concentration, topsoil lead concentration and topsoil zinc concentration. Figure 2 shows the map with locations of the sample. This dataset is characterized by the absence of zero values.
Figure 2: The flood plain of the river Meuse in Netherlands.
We have selected 3 covariates to associate the components with, namely the relative elevation above local river bed (in metres), the organic matter, kg (100 kg)-1 soil (percent) and the distance to river Meuse (in metres), as obtained during the field survey.
[…]
Table 3: Meuse river data set: average results regarding the optimal choice of $\alpha$, $k$, $h$, KLD and running time (in seconds) for each of the three models.
Model KLD $\alpha$ $k$ $h$ Running time
$\alpha$–regression 0.006 0.500 1.017
$\alpha$–SLX 0.006 0.500 3 29.868
$\alpha$–SAR 0.006 0.435 4 370.723
[…]
We then fitted the regression models using the optimal parameters obtained based on the CV protocol and computed the correlations (component-wise) between the observed and fitted compositions. Table 2 contains these correlations. Evidently, the $\alpha$–SAR model seems to have achieved the best fit. It is important to highlight that the spatial autoregressive parameter $\rho$ of the $\alpha$–SAR model was equal to -0.206 with a standard error equal to 0.164. Table 4 contains the correlations between each pair of the observed and fitted components for each of the three regression models. The $\alpha$–regression, $\alpha$-SLX and the $\alpha$–SAR model have performed equally well.
Table 4: Meuse river data set: Pearson correlations between each pair of the observed and fitted components for each of the three regression models.
Cadmium Copper Lead Zinc
$\alpha$–regression 0.638 0.543 0.471 0.628
$\alpha$–SLX 0.717 0.575 0.506 0.653
$\alpha$–SAR 0.592 0.559 0.472 0.634
[…]
5 Conclusions
[…]
The relevant functions to perform all the $\alpha$–regression models are available in the R package CompositionalSR (compositionalsr2025), which imports the package minpack.lm.
Future research could explore nonparametric spatially varying models for compositional data, as well as hybrid approaches that blend GWR with machine learning techniques for complex compositional systems.
Multivariate extensions of Moran’s I test (lin2023) have been proposed, but perhaps we could generalise the test both in arbitrary dimensions and beyond the linear association. One could substitute the Pearson correlation by the distance correlation szekely2007. Due to the constrained nature of the compositional data, the test should be applied to the residuals of a $\alpha$ type regression model and test whether the residuals and their spatial neighboring averaged residuals are dependent or not. The p-value of independence between $\mathbf{Y}_{\alpha}$ and $\mathbf{WY}_{\alpha}$ can be computed either via permutations (szekely2007) or asymptotically via the $\chi^{2}$ distribution (shen2022).
Appendix
Appendix A Univariate spatial regression models
A.1 The SLX model
The SLX model provides a useful and interpretable framework for identifying spatial spillover effects through covariates alone. While it lacks the feedback mechanisms of models that include $Wy$ (spatial autocorrelation of the dependent variable), it remains a robust and easily estimable tool for exploring spatial interactions. The structure of the SLX model allows researchers to capture how characteristics of neighboring spatial units affect local outcomes without introducing simultaneity. The general form of the SLX model is
$y_{i}=\beta_{0}+\sum_{k=1}^{p}\beta_{k}x_{ik}+\sum_{k=1}^{p}\gamma_{k}\left(\sum_{j\neq i}w_{ij}x_{ik}\right)+\varepsilon_{i},$ (A.1)
where $y$ denotes the dependent variable, $x_{k}$ denotes the kkt-h covariate, $w_{ij}$ is the $(i,j)$ element of the $n\times n$ spatial weights (contiguity) matrix $\mathbf{W}$ representing the spatial relationships between observations (e.g., contiguity or inverse distance), and $\sum_{j\neq i}w_{ij}x_{k}$ denotes the $k$-th spatially lagged covariate. The $\beta s$ and $\gamma s$ are parameters corresponding to the direct (local) and indirect (spillover) effects, respectively, and $\varepsilon$ is the classical error term.
A.2 The SAR model
The SAR model associates the response with its neighbours may be written as
$y_{i}$ $=$ $\rho\mathbf{Wy}+\mathbf{X}_{i}\mathbf{\beta}+\varepsilon_{i}$
$y_{i}$ $=$ $S(\rho)\mathbf{X}_{i}\mathbf{\beta}+S(\rho)\varepsilon_{i},$
where $S(\rho)=\left(1-\rho\mathbf{W}\right)^{-1}$ and $\rho\in(0,1)$ is the spatial autoregressive parameter and determines how much the dependent variable in one area is influenced by values in neighboring areas, according to the spatial structure defined by $\mathbf{W}$. In theory it can also take negative values, but positive values are more meaningful, in the sense that the neighbours of a farm that produces olive oil, for instance, affect the farm in a positive way.
The above model may also be written as $y_{i}=\mathbf{X}_{i}\mathbf{\gamma}+u_{i}$ and be estimated using standard linear regression. The problem though is that we cannot disentangle the components of $\mathbf{\gamma}$, we cannot identify the parameters $\beta$ and $\rho$ and for this. Secondly, due to the endogeneity caused, the estimated regression coefficients are biased. Therefore suitable techniques must be employed, unless $\rho=0$, i.e. no spatial dependence, in which case the standard linear regression suffices.
A.3 The GWR model
GWR has become a widely used technique in spatial statistics for modeling spatially varying relationships. Traditional regression assumes stationarity of relationships across space, but GWR relaxes this assumption by allowing coefficients to vary geographically (brunsdon1996). Meanwhile, compositional data–data sets where variables represent proportions of a whole and are constrained to sum to unity–have gained attention in many disciplines, including environmental sciences, geology, and social sciences. When spatial heterogeneity and compositional constraints intersect, specialized methodological developments are required. The foundational work of fotheringham2002 formalized GWR as a local regression technique that incorporates spatial weighting functions to account for the geographical location of observations.
The basic form of a standard multiple linear regression is:
$y_{i}=\beta_{0}+\sum_{k=1}^{p}\beta_{k}x_{ik}+\varepsilon_{i}.$
In GWR, the parameters are allowed to vary with location:
$y_{i}=\beta_{0}(\nu_{i},v_{i})+\sum_{k=1}^{p}\beta_{k}(\nu_{i},v_{i})\,x_{ik}+\varepsilon_{i},$
where $(\nu_{i},v_{i})$ denotes the spatial coordinates of observation $i$ ($\nu_{i}$ and $v_{i}$ typically correspond to latitude and longitude, respectively), and $\beta_{k}(\nu_{i},v_{i})$ are the location-specific parameter estimates.
For each location $(u_{i},v_{i})$, the parameter vector is estimated as:
$\widehat{\mathbf{\beta}}(\nu_{i},v_{i})=\left(X^{\top}W(\nu_{i},v_{i})X\right)^{-1}X^{\top}W(\nu_{i},v_{i})\mathbf{y},$
where $X$ is the design matrix and $W(\nu_{i},v_{i})$ is a spatial weighting matrix assigning higher weights to observations closer to $(\nu_{i},v_{i})$.
Appendix B Gradient vector and Hessian matrix for the $\alpha$–regression
The least squares objective function is
$\text{SSE}(\mathbf{Y},\mathbf{X};\alpha,\mathbf{B})=-\frac{1}{2}\text{tr}[(\mathbf{y}_{\alpha}-\mathbf{\mu}_{\alpha})^{\top}(\mathbf{y}_{\alpha}-\mathbf{\mu}_{\alpha})],$
where $\mathbf{y}_{\alpha}$ is the $\alpha$–transformed observed compositional data ($n\times d$ matrix), $\mathbf{\mu}_{\alpha}$ is the $\alpha$–transformed fitted compositional values ($n\times d$ matrix), $n$ is the number of observations, and d=D1d=D-1 where $D$ is the number of components in the composition.
The fitted compositional values come from the inverse alr transformation:
$\mu_{1}=\frac{1}{1+\sum_{j=1}^{d}e^{x^{\top}\beta_{j}}},\quad\mu_{i}=\frac{e^{x^{\top}\beta_{i-1}}}{1+\sum_{j=1}^{d}e^{x^{\top}\beta_{j}}},\quad i=2,\ldots,D.$
B.1 The $\alpha$–transformation
The $\alpha$–transformation consists of two steps:
Step 1: Power transformation
$u_{i}=\frac{\mu_{i}^{\alpha}}{\sum_{j=1}^{D}\mu_{j}^{\alpha}},\quad i=1,\ldots,D.$
Step 2: Helmert transformation
$z=\frac{1}{\alpha}H(Du-j_{D}),$
where $H$ is the $d\times D$ Helmert sub-matrix and $j_{D}$ is a $D$-dimensional vector of ones.
B.2 First Derivatives (Gradient)
B.2.1 Main Gradient Formula
$\frac{\partial l(\alpha)}{\partial\beta_{k}}=\text{tr}\left[(\mathbf{y}_{\alpha}-\mathbf{\mu}_{\alpha})^{\top}\frac{\partial\mathbf{\mu}_{\alpha}}{\partial\beta_{k}}\right].$
B.2.2 Expanded Gradient Formula
$\frac{\partial l(\alpha)}{\partial\beta_{k}}=\sum_{i=1}^{n}\sum_{m=1}^{d}\sum_{\ell=1}^{D}\sum_{p=1}^{D}r_{\alpha,im}\cdot\frac{D}{\alpha}H_{m\ell}\cdot\frac{\partial u_{i\ell}}{\partial\mu_{ip}}\cdot\frac{\partial\mu_{ip}}{\partial\beta_{k}}\cdot x_{i},$
where $r_{\alpha,im}=y_{\alpha,im}-m_{\alpha,im}$ are the residuals in $\alpha$–transformed space, $H_{m\ell}$ is the $(m,\ell)$ element of the Helmert sub-matrix, and $x_{i}$ is the covariate vector for observation $i$.
B.2.3 Jacobian of Power Transformation
$\frac{\partial u_{i\ell}}{\partial\mu_{ip}}=\begin{cases}\displaystyle\frac{\alpha\mu_{i\ell}^{\alpha-1}}{\sum_{j=1}^{D}\mu_{ij}^{\alpha}}\left(1-\frac{\mu_{i\ell}^{\alpha}}{\sum_{j=1}^{D}\mu_{ij}^{\alpha}}\right)&\text{if }\ell=p\\ [3ex]\displaystyle-\frac{\alpha\mu_{i\ell}^{\alpha}\mu_{ip}^{\alpha-1}}{(\sum_{j=1}^{D}\mu_{ij}^{\alpha})^{2}}&\text{if }\ell\neq p\end{cases}.$
Let $T_{i}=\sum_{j=1}^{D}\mu_{ij}^{\alpha}$. In compact form:
$\frac{\partial u_{i\ell}}{\partial\mu_{ip}}=\frac{\alpha\mu_{ip}^{\alpha-1}}{T_{i}}\left(\delta_{\ell p}-\frac{\mu_{i\ell}^{\alpha}}{T_{i}}\right),$
where $\delta_{\ell p}$ is the Kronecker delta.
B.2.4 Jacobian of Multinomial Logit
Let $S_{i}=1+\sum_{j=1}^{d}e^{x_{i}^{\top}\beta_{j}}$.
$\frac{\partial\mu_{ip}}{\partial\beta_{k}}=\begin{cases}-\mu_{i1}\mu_{ik}x_{i}&\text{if }p=1\\ \mu_{ik}(1-\mu_{ik})x_{i}&\text{if }p=k+1\\ -\mu_{ip}\mu_{ik}x_{i}&\text{if }p\neq 1,p\neq k+1\end{cases},$
where $\mu_{ik}=\mu_{i,k+1}$ (the $(k+1)$-th component of the composition).
B.2.5 Vectorized Gradient Formula
$\frac{\partial l(\alpha)}{\partial\beta_{k}}=X^{\top}w_{k},$
where the weight vector $w_{k}\in\mathbb{R}^{n}$ has elements:
$w_{k,i}=\left\{r_{\alpha,i}^{\top}\cdot\frac{D}{\alpha}H\cdot J_{u}(i)\cdot J_{\mu}(i,k)\right\}.$
Diagonal Contribution
$w_{k,i}^{\text{diag}}=\sum_{\ell=1}^{D}r_{\alpha,i\ell}H_{\ell}J_{u,\text{diag}}(i,\ell)J_{\mu}(i,\ell,k)$
where $J_{u,\text{diag}}(i,\ell)=\frac{\alpha\mu_{i\ell}^{\alpha-1}}{T_{i}}\left(1-\frac{\mu_{i\ell}^{\alpha}}{T_{i}}\right)$.
Off-Diagonal Contribution
$w_{k,i}^{\text{off-diag}}=-\frac{\alpha}{T_{i}^{2}}\left[\left(\sum_{\ell=1}^{D}r_{\alpha,i\ell}H_{\ell}\mu_{i\ell}^{\alpha}\right)\left(\sum_{p=1}^{D}\mu_{ip}^{\alpha-1}J_{\mu}(i,p,k)\right)-\sum_{\ell=1}^{D}r_{\alpha,i\ell}H_{\ell}\mu_{i\ell}^{\alpha}\mu_{i\ell}^{\alpha-1}J_{\mu}(i,\ell,k)\right].$
Total Weight:
$w_{k,i}=w_{k,i}^{\text{diag}}+w_{k,i}^{\text{off-diag}}.$
B.3 Hessian matrix for the $\alpha$–regression
The sum of squares of the errors is:
$l(\alpha)=-\frac{1}{2}\sum_{i=1}^{n}\sum_{j=1}^{d}(y_{\alpha,ij}-m_{\alpha,ij})^{2}.$
We will compute the Hessian matrix including all second-order terms. The gradient is
$\frac{\partial l(\alpha)}{\partial\beta_{k}}=\sum_{i=1}^{n}\sum_{j=1}^{d}r_{\alpha,ij}\frac{\partial m_{\alpha,ij}}{\partial\beta_{k}},$
where $r_{\alpha,ij}=y_{\alpha,ij}-m_{\alpha,ij}$. The structure of the Hessian matrix is:
$H_{\text{exact}}=\begin{bmatrix}H_{1,1}&H_{1,2}&\cdots&H_{1,d}\\ H_{2,1}&H_{2,2}&\cdots&H_{2,d}\\ \vdots&\vdots&\ddots&\vdots\\ H_{d,1}&H_{d,2}&\cdots&H_{d,d}\end{bmatrix}$.
Each block $H_{k,k^{\prime}}\in\mathbb{R}^{p\times p}$ includes both first and second-order terms.
The derivative with respect to $\beta_{k^{\prime}}$ is:
$\frac{\partial^{2}l(\alpha)}{\partial\beta_{k}\partial\beta_{k^{\prime}}}=\underbrace{-\sum_{i=1}^{n}\sum_{j=1}^{d}\frac{\partial m_{\alpha,ij}}{\partial\beta_{k}}\frac{\partial m_{\alpha,ij}}{\partial\beta_{k^{\prime}}}}_{\text{First-order term (GN)}}+\underbrace{\sum_{i=1}^{n}\sum_{j=1}^{d}r_{\alpha,ij}\frac{\partial^{2}m_{\alpha,ij}}{\partial\beta_{k}\partial\beta_{k^{\prime}}}}_{\text{Second-order term}}.$
B.3.1 First-Order Term (Gauss-Newton Part)
This is identical to the Gauss-Newton approximation:
$H_{k,k^{\prime}}^{(1)}=-\sum_{i=1}^{n}\sum_{j=1}^{d}\frac{\partial m_{\alpha,ij}}{\partial\beta_{k}}\frac{\partial m_{\alpha,ij}}{\partial\beta_{k^{\prime}}}=-X^{\top}\text{diag}(W_{k,k^{\prime}})X.$
where
$W_{k,k^{\prime}}(i,i)=\sum_{j=1}^{d}\frac{\partial m_{\alpha,ij}}{\partial\beta_{k}}\cdot\frac{\partial m_{\alpha,ij}}{\partial\beta_{k^{\prime}}}.$
B.3.2 Second-Order Term (Exact Correction)
Computation of $\frac{\partial^{2}m_{\alpha,ij}}{\partial\beta_{k}\partial\beta_{k^{\prime}}}$.
B.3.3 Chain Rule for Second Derivative
The chain rule for first derivative is
$\frac{\partial m_{\alpha,i}}{\partial\beta_{k}}=\frac{D}{\alpha}H\cdot J_{u}(i)\cdot J_{\mu}(i,k).$
Taking the derivative with respect to $\beta_{k^{\prime}}$:
$\frac{\partial^{2}m_{\alpha,i}}{\partial\beta_{k}\partial\beta_{k^{\prime}}}=\frac{D}{\alpha}H\cdot\left[\frac{\partial J_{u}(i)}{\partial\beta_{k^{\prime}}}\cdot J_{\mu}(i,k)+J_{u}(i)\cdot\frac{\partial J_{\mu}(i,k)}{\partial\beta_{k^{\prime}}}\right].$
B.3.4 Second Derivative of Power Transformation
We need $\frac{\partial J_{u}(i)}{\partial\beta_{k^{\prime}}}$, which involves $\frac{\partial^{2}u_{\ell}}{\partial\mu_{p}\partial\mu_{q}}$. Let $T_{i}=\sum_{j=1}^{D}\mu_{ij}^{\alpha}$.
Diagonal-Diagonal: $\ell=p=q$
$\frac{\partial^{2}u_{\ell}}{\partial\mu_{\ell}^{2}}=\frac{\alpha(\alpha-1)\mu_{\ell}^{\alpha-2}}{T}\left(1-\frac{\mu_{\ell}^{\alpha}}{T}\right)-\frac{2\alpha^{2}\mu_{\ell}^{2\alpha-2}}{T^{2}}+\frac{2\alpha^{2}\mu_{\ell}^{3\alpha-2}}{T^{3}}.$
Diagonal-Off-diagonal: $\ell=p\neq q$
$\frac{\partial^{2}u_{\ell}}{\partial\mu_{\ell}\partial\mu_{q}}=-\frac{\alpha(\alpha-1)\mu_{\ell}^{\alpha-1}\mu_{q}^{\alpha-1}}{T^{2}}\left(1-\frac{\mu_{\ell}^{\alpha}}{T}\right)-\frac{\alpha^{2}\mu_{\ell}^{\alpha}\mu_{q}^{\alpha-1}}{T^{2}}+\frac{2\alpha^{2}\mu_{\ell}^{2\alpha-1}\mu_{q}^{\alpha-1}}{T^{3}}.$
Off-diagonal-Off-diagonal: $\ell\neq p,\ell=q$
$\frac{\partial^{2}u_{\ell}}{\partial\mu_{\ell}\partial\mu_{p}}=-\frac{\alpha(\alpha-1)\mu_{\ell}^{\alpha-1}\mu_{p}^{\alpha-1}}{T^{2}}\left(1-\frac{\mu_{\ell}^{\alpha}}{T}\right)-\frac{\alpha^{2}\mu_{\ell}^{2\alpha-1}\mu_{p}^{\alpha-1}}{T^{2}}+\frac{2\alpha^{2}\mu_{\ell}^{3\alpha-1}\mu_{p}^{\alpha-1}}{T^{3}}.$
Fully Off-diagonal: $\ell\neq p,\ell\neq q,p\neq q$
$\frac{\partial^{2}u_{\ell}}{\partial\mu_{p}\partial\mu_{q}}=-\frac{\alpha(\alpha-1)\mu_{\ell}^{\alpha}\mu_{p}^{\alpha-2}\delta_{pq}}{T^{2}}-\frac{\alpha^{2}\mu_{\ell}^{\alpha}\mu_{p}^{\alpha-1}\mu_{q}^{\alpha-1}}{T^{2}}+\frac{2\alpha^{2}\mu_{\ell}^{\alpha}\mu_{p}^{\alpha-1}\mu_{q}^{\alpha-1}}{T^{3}},$
where $\delta_{pq}$ is the Kronecker delta.
B.3.5 General Formula for Hessian of Power Transformation
Let $H_{u}(i,\ell)$ denote the $D\times D$ Hessian matrix for component $\ell$ of $u_{i}$:
$H_{u}(i,\ell)]_{pq}=\frac{\partial^{2}u_{i\ell}}{\partial\mu_{ip}\partial\mu_{iq}}.$
Then:
$\frac{\partial J_{u}(i)}{\partial\beta_{k^{\prime}}}=\sum_{\ell=1}^{D}\sum_{p=1}^{D}H_{u}(i,\ell)_{pq}\cdot\frac{\partial\mu_{ip}}{\partial\beta_{k^{\prime}}}\cdot e_{\ell}e_{q}^{\top}$
where $e_{\ell}$ is the $\ell$-th standard basis vector. This becomes a $D\times D$ matrix where each element is:
$\left[\frac{\partial J_{u}(i)}{\partial\beta_{k^{\prime}}}\right]_{\ell q}=\sum_{p=1}^{D}\frac{\partial^{2}u_{i\ell}}{\partial\mu_{ip}\partial\mu_{iq}}\cdot\frac{\partial\mu_{ip}}{\partial\beta_{k^{\prime}}}.$
B.3.6 Second Derivative of Multinomial Logit
We need $\frac{\partial J_{\mu}(i,k)}{\partial\beta_{k^{\prime}}}$, which involves $\frac{\partial^{2}\mu_{ip}}{\partial\beta_{k}\partial\beta_{k^{\prime}}}$. Let $\mu_{ik}=\mu_{i,k+1}$ and $\mu_{ik^{\prime}}=\mu_{i,k^{\prime}+1}$.
For component $p=1$ (reference):
$\frac{\partial^{2}\mu_{i1}}{\partial\beta_{k}^{2}}=\mu_{i1}\mu_{ik}(\mu_{ik}-\mu_{i1})x_{i}x_{i}^{\top}.$
For component $p=k+1$:
$\frac{\partial^{2}\mu_{i,k+1}}{\partial\beta_{k}^{2}}=\mu_{ik}(1-\mu_{ik})(1-2\mu_{ik})x_{i}x_{i}^{\top}.$
For other components $p\neq 1,p\neq k+1$:
$\frac{\partial^{2}\mu_{ip}}{\partial\beta_{k}^{2}}=\mu_{ip}\mu_{ik}(\mu_{ik}+\mu_{ip})x_{i}x_{i}^{\top}.$
Case 2: $k\neq k^{\prime}$ (Different Components)
For component $p=1$ (reference):
$\frac{\partial^{2}\mu_{i1}}{\partial\beta_{k}\partial\beta_{k^{\prime}}}=\mu_{i1}\mu_{ik}\mu_{ik^{\prime}}x_{i}x_{i}^{\top}.$
For component $p=k+1$:
$\frac{\partial^{2}\mu_{i,k+1}}{\partial\beta_{k}\partial\beta_{k^{\prime}}}=-\mu_{ik}\mu_{ik^{\prime}}(1-\mu_{ik})x_{i}x_{i}^{\top}.$
For component $p=k^{\prime}+1$:
$\frac{\partial^{2}\mu_{i,k^{\prime}+1}}{\partial\beta_{k}\partial\beta_{k^{\prime}}}=-\mu_{ik^{\prime}}\mu_{ik}(1-\mu_{ik^{\prime}})x_{i}x_{i}^{\top}.$
For other components $p\neq 1,k+1,k^{\prime}+1$:
$\frac{\partial^{2}\mu_{ip}}{\partial\beta_{k}\partial\beta_{k^{\prime}}}=\mu_{ip}\mu_{ik}\mu_{ik^{\prime}}x_{i}x_{i}^{\top}.$
B.3.7 Assembling the Second-Order Term
The second-order correction to the Hessian is:
$H_{k,k^{\prime}}^{(2)}=\sum_{i=1}^{n}\sum_{j=1}^{d}r_{\alpha,ij}\frac{\partial^{2}m_{\alpha,ij}}{\partial\beta_{k}\partial\beta_{k^{\prime}}},$
where
$\frac{\partial^{2}m_{\alpha,ij}}{\partial\beta_{k}\partial\beta_{k^{\prime}}}=[H]_{j}\cdot\left[\frac{\partial J_{u}(i)}{\partial\beta_{k^{\prime}}}\cdot J_{\mu}(i,k)+J_{u}(i)\cdot\frac{\partial J_{\mu}(i,k)}{\partial\beta_{k^{\prime}}}\right]\cdot\frac{D}{\alpha}.$
Here $H]_{j}$ denotes the $j$-th row of the Helmert matrix.
Explicit Form
$H_{k,k^{\prime}}^{(2)}=\sum_{i=1}^{n}r_{\alpha,i}^{\top}\cdot\frac{D}{\alpha}H\cdot\left[\sum_{p=1}^{D}\left(\sum_{\ell,q}H_{u}(i,\ell,p,q)\frac{\partial\mu_{ip}}{\partial\beta_{k^{\prime}}}\right)J_{\mu}(i,k)_{q}e_{\ell}e_{q}^{\top}+J_{u}(i)\cdot\frac{\partial^{2}\mu}{\partial\beta_{k}\partial\beta_{k^{\prime}}}x_{i}x_{i}^{\top}\right].$
B.3.8 Complete Hessian (Exact)
$H_{k,k^{\prime}}=H_{k,k^{\prime}}^{(1)}+H_{k,k^{\prime}}^{(2)}=-X^{\top}\text{diag}(W_{k,k^{\prime}}^{(1)})X+\sum_{i=1}^{n}r_{\alpha,i}^{\top}\cdot S_{k,k^{\prime}}(i)\cdot x_{i}x_{i}^{\top},$
where $W_{k,k^{\prime}}^{(1)}$ are the Gauss-Newton weights and $S_{k,k^{\prime}}(i)$ is the second-order correction tensor for observation $i$.
B.4 Gradient vecctor and Hessian matrix for the $\alpha$–SAR model
The $\alpha$–SAR model minimizes the sum of squared errors (SSE):
$\text{SSE}(\mathbf{Y},\mathbf{X};\alpha,\rho,\mathbf{B})=\ell\left(\mathbf{\theta}\right)=\sum_{i=1}^{n}\|\mathbf{y}_{i,\alpha}-\mathbf{\mu}_{i,\alpha}\|^{2}=\sum_{i=1}^{n}(\mathbf{y}_{i,\alpha}-\mathbf{\mu}_{i,\alpha})^{\top}(\mathbf{y}_{i,\alpha}-\mathbf{\mu}_{i,\alpha})$
where $\mathbf{\theta}=(\mathop{\operator@font vec}\nolimits(\mathbf{B})^{\top},\rho)^{\top}$ contains all parameters.
Appendix C Model Specification Review
The fitted compositional values are:
$\mu_{i}=\begin{cases}\dfrac{1}{1+\sum\limits_{j=1}^{d}e^{\tilde{\mathbf{x}}_{i}^{\top}\mathbf{\beta}_{j}}}&\text{for }i=1\\[8.61108pt] \dfrac{e^{\tilde{\mathbf{x}}_{i}^{\top}\mathbf{\beta}_{i-1}}}{1+\sum\limits_{j=1}^{d}e^{\tilde{\mathbf{x}}_{i}^{\top}\mathbf{\beta}_{j}}}&\text{for }i=2,\ldots,D\end{cases}$
where:
$\tilde{\mathbf{x}}_{i}=[\mathbf{S}(\rho)^{-1}\mathbf{x}]_{i}=[(I_{n}-\rho W)^{-1}\mathbf{x}]_{i}$
The Hessian matrix has the block structure:
$\mathbf{H}=\begin{bmatrix}\mathbf{H}_{\mathbf{\beta}\mathbf{\beta}}&\mathbf{H}_{\mathbf{\beta}\rho}\\ \mathbf{H}_{\rho\mathbf{\beta}}&H_{\rho\rho}\end{bmatrix}$,
where:
  • $\mathbf{H}_{\mathbf{\beta}\mathbf{\beta}}$ is $(dp)\times(dp)$: derivatives w.r.t. regression coefficients
  • $\mathbf{H}_{\mathbf{\beta}\rho}$ is $(dp)\times 1$: mixed derivatives
  • $\mathbf{H}_{\rho\mathbf{\beta}}=\mathbf{H}_{\mathbf{\beta}\rho}^{\top}$ by symmetry
  • $H_{\rho\rho}$ is $1\times 1$: second derivative w.r.t. $\rho$
C.1 First Derivatives (Gradient)
C.1.1 Gradient with respect to $\mathbf{\beta}_{k}$
$\frac{\partial\ell}{\partial\mathbf{\beta}_{k}}=\sum_{i=1}^{n}\sum_{j=1}^{d}r_{i,\alpha,j}\frac{\partial\mu_{i,\alpha,j}}{\partial\mathbf{\beta}_{k}},$
where $r_{i,\alpha,j}=y_{i,\alpha,j}-\mu_{i,\alpha,j}$ are the residuals.
Using the chain rule:
$\frac{\partial\mu_{i,\alpha,j}}{\partial\mathbf{\beta}_{k}}=\frac{D}{\alpha}\sum_{\ell=1}^{D}\sum_{p=1}^{D}H_{j\ell}\frac{\partial u_{i\ell}}{\partial\mu_{ip}}\frac{\partial\mu_{ip}}{\partial\mathbf{\beta}_{k}},$
where $H_{j\ell}$ is the $(j,\ell)$ element of the Helmert sub-matrix.
C.1.2 Gradient with respect to $\rho$
$\frac{\partial\ell}{\partial\rho}=\sum_{i=1}^{n}\sum_{j=1}^{d}r_{i,\alpha,j}\frac{\partial\mu_{i,\alpha,j}}{\partial\rho}.$
Using the chain rule through $\tilde{\mathbf{x}}$:
$\frac{\partial\mu_{i,\alpha,j}}{\partial\rho}=\sum_{p=1}^{D}\frac{\partial\mu_{i,\alpha,j}}{\partial\mu_{ip}}\frac{\partial\mu_{ip}}{\partial\rho},$
where:
$\frac{\partial\mu_{ip}}{\partial\rho}=\sum_{s=1}^{p}\frac{\partial\mu_{ip}}{\partial\tilde{x}_{is}}\frac{\partial\tilde{x}_{is}}{\partial\rho}.$
C.2 Second Derivatives: $\mathbf{H}_{\mathbf{\beta}\mathbf{\beta}}$
The $(k,k^{\prime})$ block of the Hessian (w.r.t. $\mathbf{\beta}_{k}$ and $\mathbf{\beta}_{k^{\prime}}$) is:
$\frac{\partial^{2}\ell}{\partial\mathbf{\beta}_{k}\partial\mathbf{\beta}_{k^{\prime}}^{\top}}=-\sum_{i=1}^{n}\sum_{j=1}^{d}\frac{\partial\mu_{i,\alpha,j}}{\partial\mathbf{\beta}_{k}}\frac{\partial\mu_{i,\alpha,j}}{\partial\mathbf{\beta}_{k^{\prime}}^{\top}}+\sum_{i=1}^{n}\sum_{j=1}^{d}r_{i,\alpha,j}\frac{\partial^{2}\mu_{i,\alpha,j}}{\partial\mathbf{\beta}_{k}\partial\mathbf{\beta}_{k^{\prime}}^{\top}}.$
C.2.1 Gauss-Newton Approximation (First Term)
The first term is the Gauss-Newton approximation:
$\mathbf{H}^{(1)}_{k,k^{\prime}}=-\sum_{i=1}^{n}\sum_{j=1}^{d}\frac{\partial\mu_{i,\alpha,j}}{\partial\mathbf{\beta}_{k}}\frac{\partial\mu_{i,\alpha,j}}{\partial\mathbf{\beta}_{k^{\prime}}^{\top}}.$
This can be written in matrix form:
$\mathbf{H}^{(1)}_{k,k^{\prime}}=-\tilde{\mathbf{X}}^{\top}\operatorname{diag}(\mathbf{W}_{k,k^{\prime}})\tilde{\mathbf{X}}$,
where $\tilde{\mathbf{X}}=\mathbf{S}(\rho)^{-1}\mathbf{X}$ and:
$W_{k,k^{\prime}}(i,i)=\sum_{j=1}^{d}\frac{\partial\mu_{i,\alpha,j}}{\partial\mathbf{\beta}_{k}}\cdot\frac{\partial\mu_{i,\alpha,j}}{\partial\mathbf{\beta}_{k^{\prime}}^{\top}}.$
C.2.2 Exact Correction (Second Term)
The second term involves second derivatives of the transformed composition:
$\mathbf{H}^{(2)}_{k,k^{\prime}}=\sum_{i=1}^{n}\sum_{j=1}^{d}r_{i,\alpha,j}\frac{\partial^{2}\mu_{i,\alpha,j}}{\partial\mathbf{\beta}_{k}\partial\mathbf{\beta}_{k^{\prime}}^{\top}}.$
Computing $\frac{\partial^{2}\mu_{i,\alpha,j}}{\partial\mathbf{\beta}_{k}\partial\mathbf{\beta}_{k^{\prime}}^{\top}}$ requires:
$\frac{\partial^{2}\mu_{i,\alpha,j}}{\partial\mathbf{\beta}_{k}\partial\mathbf{\beta}_{k^{\prime}}^{\top}}=\frac{D}{\alpha}\sum_{\ell=1}^{D}\sum_{p=1}^{D}H_{j\ell}\left[\frac{\partial^{2}u_{i\ell}}{\partial\mu_{ip}\partial\mathbf{\beta}_{k^{\prime}}^{\top}}\frac{\partial\mu_{ip}}{\partial\mathbf{\beta}_{k}}+\frac{\partial u_{i\ell}}{\partial\mu_{ip}}\frac{\partial^{2}\mu_{ip}}{\partial\mathbf{\beta}_{k}\partial\mathbf{\beta}_{k^{\prime}}^{\top}}\right].$
C.2.3 Second Derivative of Power Transformation
$\frac{\partial^{2}u_{i\ell}}{\partial\mu_{ip}\partial\mathbf{\beta}_{k^{\prime}}^{\top}}=\sum_{q=1}^{D}\frac{\partial^{2}u_{i\ell}}{\partial\mu_{ip}\partial\mu_{iq}}\frac{\partial\mu_{iq}}{\partial\mathbf{\beta}_{k^{\prime}}^{\top}}.$
The second derivatives of the power transformation $\frac{\partial^{2}u_{i\ell}}{\partial\mu_{ip}\partial\mu_{iq}}$ are provided in Appendix A of the original paper.
C.2.4 Second Derivative of Multinomial Logit
For the multinomial logit part, the second derivative depends on the case:
Case 1: $k=k^{\prime}$ (same component)
For component $p=1$ (reference):
$\frac{\partial^{2}\mu_{i1}}{\partial\mathbf{\beta}_{k}^{2}}=\mu_{i1}\mu_{ik+1}(\mu_{ik+1}-\mu_{i1})\tilde{\mathbf{x}}_{i}\tilde{\mathbf{x}}_{i}^{\top}.$
For component $p=k+1$:
$\frac{\partial^{2}\mu_{i,k+1}}{\partial\mathbf{\beta}_{k}^{2}}=\mu_{ik+1}(1-\mu_{ik+1})(1-2\mu_{ik+1})\tilde{\mathbf{x}}_{i}\tilde{\mathbf{x}}_{i}^{\top}.$
For other components $p\neq 1,p\neq k+1$:
$\frac{\partial^{2}\mu_{ip}}{\partial\mathbf{\beta}_{k}^{2}}=\mu_{ip}\mu_{ik+1}(\mu_{ik+1}+\mu_{ip})\tilde{\mathbf{x}}_{i}\tilde{\mathbf{x}}_{i}^{\top}.$
Case 2: $k\neq k^{\prime}$ (different components)
For component $p=1$ (reference):
$\frac{\partial^{2}\mu_{i1}}{\partial\mathbf{\beta}_{k}\partial\mathbf{\beta}_{k^{\prime}}^{\top}}=\mu_{i1}\mu_{ik+1}\mu_{ik^{\prime}+1}\tilde{\mathbf{x}}_{i}\tilde{\mathbf{x}}_{i}^{\top}.$
For component $p=k+1$:
$\frac{\partial^{2}\mu_{i,k+1}}{\partial\mathbf{\beta}_{k}\partial\mathbf{\beta}_{k^{\prime}}^{\top}}=-\mu_{ik+1}\mu_{ik^{\prime}+1}(1-\mu_{ik+1})\tilde{\mathbf{x}}_{i}\tilde{\mathbf{x}}_{i}^{\top}.$
For component $p=k^{\prime}+1$:
$\frac{\partial^{2}\mu_{i,k^{\prime}+1}}{\partial\mathbf{\beta}_{k}\partial\mathbf{\beta}_{k^{\prime}}^{\top}}=-\mu_{ik^{\prime}+1}\mu_{ik+1}(1-\mu_{ik^{\prime}+1})\tilde{\mathbf{x}}_{i}\tilde{\mathbf{x}}_{i}^{\top}.$
For other components $p\neq 1,k+1,k^{\prime}+1$:
$\frac{\partial^{2}\mu_{ip}}{\partial\mathbf{\beta}_{k}\partial\mathbf{\beta}_{k^{\prime}}^{\top}}=\mu_{ip}\mu_{ik+1}\mu_{ik^{\prime}+1}\tilde{\mathbf{x}}_{i}\tilde{\mathbf{x}}_{i}^{\top}.$
C.2.5 Mixed Derivatives: $\mathbf{H}_{\mathbf{\beta}\rho}$
The mixed derivative block is:
$\frac{\partial^{2}\ell}{\partial\mathbf{\beta}_{k}\partial\rho}=-\sum_{i=1}^{n}\sum_{j=1}^{d}\frac{\partial\mu_{i,\alpha,j}}{\partial\mathbf{\beta}_{k}}\frac{\partial\mu_{i,\alpha,j}}{\partial\rho}+\sum_{i=1}^{n}\sum_{j=1}^{d}r_{i,\alpha,j}\frac{\partial^{2}\mu_{i,\alpha,j}}{\partial\mathbf{\beta}_{k}\partial\rho}.$
C.2.6 Gauss-Newton Part
$\mathbf{H}^{(1)}_{\mathbf{\beta}_{k},\rho}=-\sum_{i=1}^{n}\sum_{j=1}^{d}\frac{\partial\mu_{i,\alpha,j}}{\partial\mathbf{\beta}_{k}}\frac{\partial\mu_{i,\alpha,j}}{\partial\rho}.$
C.2.7 Exact Correction
$\mathbf{H}^{(2)}_{\mathbf{\beta}_{k},\rho}=\sum_{i=1}^{n}\sum_{j=1}^{d}r_{i,\alpha,j}\frac{\partial^{2}\mu_{i,\alpha,j}}{\partial\mathbf{\beta}_{k}\partial\rho}.$
The mixed derivative requires:
$\frac{\partial^{2}\mu_{i,\alpha,j}}{\partial\mathbf{\beta}_{k}\partial\rho}=\frac{D}{\alpha}\sum_{\ell=1}^{D}\sum_{p=1}^{D}H_{j\ell}\left[\frac{\partial^{2}u_{i\ell}}{\partial\mu_{ip}\partial\rho}\frac{\partial\mu_{ip}}{\partial\mathbf{\beta}_{k}}+\frac{\partial u_{i\ell}}{\partial\mu_{ip}}\frac{\partial^{2}\mu_{ip}}{\partial\mathbf{\beta}_{k}\partial\rho}\right].$
C.2.8 Computing $\frac{\partial^{2}\mu_{ip}}{\partial\mathbf{\beta}_{k}\partial\rho}$
Using the chain rule:
$\frac{\partial^{2}\mu_{ip}}{\partial\mathbf{\beta}_{k}\partial\rho}=\sum_{q=1}^{D}\frac{\partial^{2}\mu_{ip}}{\partial\mu_{iq}\partial\mathbf{\beta}_{k}}\frac{\partial\mu_{iq}}{\partial\rho}+\frac{\partial\mu_{ip}}{\partial\mathbf{\beta}_{k}}\frac{\partial\tilde{\mathbf{x}}_{i}}{\partial\rho}.$
The key term is:
$\frac{\partial}{\partial\rho}\left[\frac{\partial\mu_{ip}}{\partial\mathbf{\beta}_{k}}\right]=\frac{\partial\mu_{ip}}{\partial\mathbf{\beta}_{k}\partial\tilde{\mathbf{x}}_{i}}\frac{\partial\tilde{\mathbf{x}}_{i}}{\partial\rho}.$
Since $\frac{\partial\tilde{\mathbf{x}}_{i}}{\partial\rho}=\mathbf{S}(\rho)^{-1}W\tilde{\mathbf{x}}_{i}$, we have the following expressions.
For component $p=1$:
$\frac{\partial^{2}\mu_{i1}}{\partial\mathbf{\beta}_{k}\partial\rho}=-\mu_{i1}\mu_{ik+1}[\mathbf{S}(\rho)^{-1}W\tilde{\mathbf{x}}_{i}]+\mu_{i1}\mu_{ik+1}(2\mu_{i1}-1)\tilde{\mathbf{x}}_{i}\cdot[\mathbf{S}(\rho)^{-1}W\tilde{\mathbf{x}}_{i}]^{\top}\mathbf{\beta}_{k}.$
For component $p=k+1$:
$\frac{\partial^{2}\mu_{i,k+1}}{\partial\mathbf{\beta}_{k}\partial\rho}=\mu_{ik+1}(1-\mu_{ik+1})[\mathbf{S}(\rho)^{-1}W\tilde{\mathbf{x}}_{i}]+\mu_{ik+1}(1-\mu_{ik+1})(1-2\mu_{ik+1})\tilde{\mathbf{x}}_{i}\cdot[\mathbf{S}(\rho)^{-1}W\tilde{\mathbf{x}}_{i}]^{\top}\mathbf{\beta}_{k}.$
For other components $p\neq 1,k+1$:
$\frac{\partial^{2}\mu_{ip}}{\partial\mathbf{\beta}_{k}\partial\rho}=-\mu_{ip}\mu_{ik+1}[\mathbf{S}(\rho)^{-1}W\tilde{\mathbf{x}}_{i}]+\mu_{ip}\mu_{ik+1}(\mu_{ip}+\mu_{ik+1})\tilde{\mathbf{x}}_{i}\cdot[\mathbf{S}(\rho)^{-1}W\tilde{\mathbf{x}}_{i}]^{\top}\mathbf{\beta}_{k}.$
C.2.9 Second Derivative: $H_{\rho\rho}$
The pure second derivative with respect to $\rho$ is:
$\frac{\partial^{2}\ell}{\partial\rho^{2}}=-\sum_{i=1}^{n}\sum_{j=1}^{d}\left(\frac{\partial\mu_{i,\alpha,j}}{\partial\rho}\right)^{2}+\sum_{i=1}^{n}\sum_{j=1}^{d}r_{i,\alpha,j}\frac{\partial^{2}\mu_{i,\alpha,j}}{\partial\rho^{2}}.$
C.2.10 Gauss-Newton Part
$H^{(1)}_{\rho\rho}=-\sum_{i=1}^{n}\sum_{j=1}^{d}\left(\frac{\partial\mu_{i,\alpha,j}}{\partial\rho}\right)^{2}.$
C.2.11 Exact Correction
$H^{(2)}_{\rho\rho}=\sum_{i=1}^{n}\sum_{j=1}^{d}r_{i,\alpha,j}\frac{\partial^{2}\mu_{i,\alpha,j}}{\partial\rho^{2}}.$
The second derivative requires:
$\frac{\partial^{2}\mu_{i,\alpha,j}}{\partial\rho^{2}}=\sum_{p=1}^{D}\frac{\partial^{2}\mu_{i,\alpha,j}}{\partial\mu_{ip}^{2}}\left(\frac{\partial\mu_{ip}}{\partial\rho}\right)^{2}+\sum_{p=1}^{D}\frac{\partial\mu_{i,\alpha,j}}{\partial\mu_{ip}}\frac{\partial^{2}\mu_{ip}}{\partial\rho^{2}}.$
C.2.12 Computing $\frac{\partial^{2}\mu_{ip}}{\partial\rho^{2}}$
This requires the second derivative of $\tilde{\mathbf{x}}_{i}$ with respect to $\rho$:
$\frac{\partial^{2}\tilde{\mathbf{x}}_{i}}{\partial\rho^{2}}$ $=$ $\frac{\partial}{\partial\rho}[\mathbf{S}(\rho)^{-1}W\tilde{\mathbf{x}}_{i}]=\mathbf{S}(\rho)^{-1}W\mathbf{S}(\rho)^{-1}W\tilde{\mathbf{x}}_{i}+\mathbf{S}(\rho)^{-1}W\frac{\partial\tilde{\mathbf{x}}_{i}}{\partial\rho}$
$\frac{\partial^{2}\tilde{\mathbf{x}}_{i}}{\partial\rho^{2}}$ $=$ $2\mathbf{S}(\rho)^{-1}W\mathbf{S}(\rho)^{-1}W\tilde{\mathbf{x}}_{i}.$
Then:
$\frac{\partial^{2}\mu_{ip}}{\partial\rho^{2}}=\sum_{q=1}^{D}\frac{\partial^{2}\mu_{ip}}{\partial\mu_{iq}^{2}}\left(\frac{\partial\mu_{iq}}{\partial\rho}\right)^{2}+\sum_{q=1}^{D}\frac{\partial\mu_{ip}}{\partial\mu_{iq}}\frac{\partial^{2}\mu_{iq}}{\partial\rho^{2}}+\sum_{s=1}^{p}\frac{\partial\mu_{ip}}{\partial\tilde{x}_{is}}\frac{\partial^{2}\tilde{x}_{is}}{\partial\rho^{2}}.$
For component $p=1$:
$\frac{\partial^{2}\mu_{i1}}{\partial\rho^{2}}=\mu_{i1}\sum_{k=1}^{d}\mu_{ik+1}\left[(\mathbf{S}(\rho)^{-1}W\tilde{\mathbf{x}}_{i})^{\top}\mathbf{\beta}_{k}\right]^{2}(2\mu_{i1}-1)-\mu_{i1}\sum_{k=1}^{d}\mu_{ik+1}\cdot 2[\mathbf{S}(\rho)^{-1}W\mathbf{S}(\rho)^{-1}W\tilde{\mathbf{x}}_{i}]^{\top}\mathbf{\beta}_{k}.$
For component $p=k+1$:
$\frac{\partial^{2}\mu_{ik+1}}{\partial\rho^{2}}=\mu_{ik+1}(1-\mu_{ik+1})[(\mathbf{S}(\rho)^{-1}W\tilde{\mathbf{x}}_{i})^{\top}\mathbf{\beta}_{k}]^{2}(1-2\mu_{ik+1})+\mu_{ik+1}(1-\mu_{ik+1})\cdot 2[\mathbf{S}(\rho)^{-1}W\mathbf{S}(\rho)^{-1}W\tilde{\mathbf{x}}_{i}]^{\top}\mathbf{\beta}_{k}.$
For other components $p\neq 1,k+1$:
$\frac{\partial^{2}\mu_{ip}}{\partial\rho^{2}}=\mu_{ip}\sum_{k=1}^{d}\mu_{ik+1}[(\mathbf{S}(\rho)^{-1}W\tilde{\mathbf{x}}_{i})^{\top}\mathbf{\beta}_{k}]^{2}(\mu_{ip}+\mu_{ik+1})-\mu_{ip}\sum_{k=1}^{d}\mu_{ik+1}\cdot 2[\mathbf{S}(\rho)^{-1}W\mathbf{S}(\rho)^{-1}W\tilde{\mathbf{x}}_{i}]^{\top}\mathbf{\beta}_{k}.$
C.2.13 Complete Hessian Matrix
The complete Hessian is:
$\mathbf{H}=\mathbf{H}^{(1)}+\mathbf{H}^{(2)},$
where $\mathbf{H}^{(1)}$ is the Gauss-Newton approximation (always negative semi-definite) and $\mathbf{H}^{(2)}$ is the exact correction involving second derivatives weighted by residuals.
C.2.14 Simplified Structure
In matrix-vector notation:
$\mathbf{H}=\begin{bmatrix}-\tilde{\mathbf{X}}^{\top}\mathbf{W}_{\mathbf{\beta}\mathbf{\beta}}\tilde{\mathbf{X}}+\mathbf{R}_{\mathbf{\beta}\mathbf{\beta}}&\mathbf{h}_{\mathbf{\beta}\rho}\\ \mathbf{h}_{\mathbf{\beta}\rho}^{\top}&h_{\rho\rho}\end{bmatrix}$,
where:
  • $\mathbf{W}_{\mathbf{\beta}\mathbf{\beta}}$ contains the Gauss-Newton weights
  • $\mathbf{R}_{\mathbf{\beta}\mathbf{\beta}}$ contains the residual-weighted second derivatives
  • $\mathbf{h}_{\mathbf{\beta}\rho}$ contains the mixed derivatives
  • $h_{\rho\rho}$ is the second derivative w.r.t. $\rho$
Appendix D Standard error of the SMEs of the $\alpha$–SLX model
The standard errors of the SMEs are computed again via the delta method, similarly to (12). Thus we need to compute the Jacobians. For observation $i$, component $\ell$, and covariate $k$ we can define the following cases:
Jacobian for the direct SMEs
$\frac{\partial DE_{i\ell k}}{\partial\beta_{ms}}=\begin{cases}\displaystyle-\delta_{sk}\mu_{i1}\mu_{im+1}-\frac{\partial\mu_{i1}}{\partial\beta_{ms}}\sum_{j=1}^{d}\beta_{jk}\mu_{ij+1}-\mu_{i1}\sum_{j=1}^{d}\beta_{jk}\frac{\partial\mu_{ij+1}}{\partial\beta_{ms}}&\text{for }\ell=1\\ \displaystyle\delta_{sk}\delta_{m,\ell-1}\mu_{i\ell}-\delta_{sk}\mu_{i\ell}\mu_{im+1}+\frac{\partial\mu_{i\ell}}{\partial\beta_{ms}}\left(\beta_{\ell-1,k}-\sum_{j=1}^{d}\beta_{jk}\mu_{ij+1}\right)\\ -\mu_{i\ell}\sum_{j=1}^{d}\beta_{jk}\frac{\partial\mu_{ij+1}}{\partial\beta_{ms}}&\text{for }\ell=2,\ldots,D.\end{cases}$
Jacobian for the indirect SMEs
$\frac{\partial IDE_{i\ell k}}{\partial\gamma_{ms}}=\begin{cases}\displaystyle-\delta_{sk}\mu_{i1}\mu_{im+1}-\frac{\partial\mu_{i1}}{\partial\gamma_{ms}}\sum_{j=1}^{d}\gamma_{jk}\mu_{ij+1}-\mu_{i1}\sum_{j=1}^{d}\gamma_{jk}\frac{\partial\mu_{ij+1}}{\partial\gamma_{ms}}&\text{for }\ell=1\\[12.91663pt] \displaystyle\delta_{sk}\delta_{m,\ell-1}\mu_{i\ell}-\delta_{sk}\mu_{i\ell}\mu_{im+1}+\frac{\partial\mu_{i\ell}}{\partial\gamma_{ms}}\left(\gamma_{\ell-1,k}-\sum_{j=1}^{d}\gamma_{jk}\mu_{ij+1}\right)&\\ \displaystyle-\mu_{i\ell}\sum_{j=1}^{d}\gamma_{jk}\frac{\partial\mu_{ij+1}}{\partial\gamma_{ms}}&\text{for }\ell=2,\ldots,D.\end{cases}$
For both cases, the necessary derivatives of $\mu$ with respect to each $\beta$ and $\gamma$ coefficient are given by
$\frac{\partial\mu_{ip}}{\partial\beta_{rs}}=\begin{cases}-\mu_{i1}\mu_{ir+1}x_{is}&\text{if }p=1\\[4.30554pt] \mu_{ip}(1-\mu_{ip})x_{is}&\text{if }p=r+1\\[4.30554pt] -\mu_{ip}\mu_{ir+1}x_{is}&\text{otherwise}\end{cases}\ \ \text{and}\ \ \frac{\partial\mu_{ip}}{\partial\gamma_{rs}}=\begin{cases}-\mu_{i1}\mu_{ir+1}(\mathbf{W}\mathbf{x})_{is}&\text{if }p=1\\[4.30554pt] \mu_{ip}(1-\mu_{ip})(\mathbf{W}\mathbf{x})_{is}&\text{if }p=r+1\\[4.30554pt] -\mu_{ip}\mu_{ir+1}(\mathbf{W}\mathbf{x})_{is}&\text{otherwise}.\end{cases}$
The Jacobian for the total SMEs is simply the sum of the Jacobians for the direct and indirect SMEs. The covariance matrices for each type of SME is the same as in the case of the $\alpha$–regression, Eqs. (12) and (13).
Appendix E Standard error of the SMEs of the $\alpha$–SAR model
For the SMEs, we need to account for uncertainty in the parameters $\mathbf{\theta}=(\text{vec}(\mathbf{B})^{\top},\rho)^{\top}$, where $\mathbf{B}=(\mathbf{\beta}_{1},\ldots,\mathbf{\beta}_{d})$ is the $d\times p$ matrix of coefficients. The complete Jacobian for observation $i$, component $\ell$, covariate $k$ has dimension $1\times(dp+1)$:
$\mathbf{J}_{i\ell k}=\left[\frac{\partial ME_{i\ell k}}{\partial\beta_{11}},\ldots,\frac{\partial ME_{i\ell k}}{\partial\beta_{dp}},\frac{\partial ME_{i\ell k}}{\partial\rho}\right]$
The effective covariate is
$\tilde{x}=S(\rho)^{-1}x=(I_{n}-\rho W)^{-1}x.$
Jacobian of the direct SMEs
The derivative of the direct SME is
$\frac{\partial DE_{i\ell k}}{\partial\beta_{rs}}=\frac{\partial^{2}\mu_{i\ell}}{\partial\tilde{x}_{k}\partial\beta_{rs}}\cdot[S(\rho)^{-1}]_{ii}$
This requires the second derivative of the composition with respect to the transformed covariate and the regression coefficients. From the multinomial logit structure:
$\frac{\partial\mu_{ip}}{\partial\beta_{rs}}=\begin{cases}-\mu_{i1}\mu_{ir+1}\tilde{x}_{is}&\text{if }p=1\\[4.30554pt] \mu_{ip}(1-\mu_{ip})\tilde{x}_{is}&\text{if }p=r+1\\[4.30554pt] -\mu_{ip}\mu_{ir+1}\tilde{x}_{is}&\text{otherwise},\end{cases}$
where $\tilde{x}_{is}$ is the $s$-th component of $\tilde{x}_{i}=S(\rho)^{-1}x_{i}$.
For $\ell=1$ (reference category):
$\frac{\partial^{2}\mu_{i1}}{\partial\tilde{x}_{k}\partial\beta_{rs}}=-\frac{\partial\mu_{i1}}{\partial\beta_{rs}}\sum_{j=1}^{d}\beta_{jk}\mu_{ij+1}-\mu_{i1}\sum_{j=1}^{d}\beta_{jk}\frac{\partial\mu_{ij+1}}{\partial\beta_{rs}}-\mu_{i1}\delta_{rk}\mu_{ir+1},$
where $\delta_{rk}$ is the Kronecker delta ($\delta_{rk}=1$ if $r=k$, 0 otherwise). For $\ell=2,\ldots,D$:
$\frac{\partial^{2}\mu_{i\ell}}{\partial\tilde{x}_{k}\partial\beta_{rs}}=\frac{\partial\mu_{i\ell}}{\partial\beta_{rs}}\left[\beta_{\ell-1,k}-\sum_{j=1}^{d}\beta_{jk}\mu_{ij+1}\right]+\mu_{i\ell}\left[\delta_{\ell-1,r}\delta_{sk}-\delta_{rk}\mu_{ir+1}-\sum_{j=1}^{d}\beta_{jk}\frac{\partial\mu_{ij+1}}{\partial\beta_{rs}}\right]$
The complete formula for $\frac{\partial DE_{i\ell k}}{\partial\beta_{rs}}$ is
For $\ell=1$:
$\frac{\partial DE_{i\ell k}}{\partial\beta_{rs}}=[S(\rho)^{-1}]_{ii}\Bigg[-\frac{\partial\mu_{i1}}{\partial\beta_{rs}}\sum_{j=1}^{d}\beta_{jk}\mu_{ij+1}-\mu_{i1}\sum_{j=1}^{d}\beta_{jk}\frac{\partial\mu_{ij+1}}{\partial\beta_{rs}}-\mu_{i1}\delta_{rk}\mu_{ir+1}\Bigg]$
For $\ell=2,\ldots,D$:
$\frac{\partial DE_{i\ell k}}{\partial\beta_{rs}}=[S(\rho)^{-1}]_{ii}\Bigg[\frac{\partial\mu_{i\ell}}{\partial\beta_{rs}}\left[\beta_{\ell-1,k}-\sum_{j=1}^{d}\beta_{jk}\mu_{ij+1}\right]+\mu_{i\ell}\left[\delta_{\ell-1,r}\delta_{sk}-\delta_{rk}\mu_{ir+1}-\sum_{j=1}^{d}\beta_{jk}\frac{\partial\mu_{ij+1}}{\partial\beta_{rs}}\right]\Bigg]$
The derivative of the direct SMEs with respect to $\rho$ involves both the change in $\mu_{i\ell}$ through $\tilde{x}$ and the change in $S(\rho)^{-1}]_{ii}$:
$\frac{\partial DE_{i\ell k}}{\partial\rho}=\frac{\partial}{\partial\rho}\left[\frac{\partial\mu_{i\ell}}{\partial\tilde{x}_{k}}\right]\cdot[S(\rho)^{-1}]_{ii}+\frac{\partial\mu_{i\ell}}{\partial\tilde{x}_{k}}\cdot\frac{\partial[S(\rho)^{-1}]_{ii}}{\partial\rho},$
where
$\frac{\partial[S(\rho)^{-1}]_{ii}}{\partial\rho}=[S(\rho)^{-1}WS(\rho)^{-1}]_{ii}$
The derivative of $\tilde{x}$ with respect to $\rho$ is
$\frac{\partial\tilde{x}}{\partial\rho}=\frac{\partial S(\rho)^{-1}}{\partial\rho}x=S(\rho)^{-1}WS(\rho)^{-1}x=S(\rho)^{-1}W\tilde{x}$
and via the chain rule application we obtain
$\frac{\partial}{\partial\rho}\left[\frac{\partial\mu_{i\ell}}{\partial\tilde{x}_{k}}\right]=\sum_{p=1}^{D}\frac{\partial^{2}\mu_{i\ell}}{\partial\tilde{x}_{k}\partial\mu_{ip}}\cdot\frac{\partial\mu_{ip}}{\partial\rho},$
where
$\frac{\partial\mu_{ip}}{\partial\rho}=\sum_{s=1}^{p}\frac{\partial\mu_{ip}}{\partial\tilde{x}_{s}}\cdot[S(\rho)^{-1}W\tilde{x}]_{is}$
Finally, we get
$\frac{\partial DE_{i\ell k}}{\partial\rho}=\left[\sum_{p=1}^{D}\frac{\partial^{2}\mu_{i\ell}}{\partial\tilde{x}_{k}\partial\mu_{ip}}\cdot\sum_{s=1}^{p}\frac{\partial\mu_{ip}}{\partial\tilde{x}_{s}}\cdot[S(\rho)^{-1}W\tilde{x}]_{is}\right]\cdot[S(\rho)^{-1}]_{ii}+\frac{\partial\mu_{i\ell}}{\partial\tilde{x}_{k}}\cdot[S(\rho)^{-1}WS(\rho)^{-1}]_{ii}.$
Jacobian for the indirect SMEs
The derivative with respect to $\beta_{rs}$ is
$\frac{\partial IE_{i\ell k}}{\partial\beta_{rs}}=\frac{\partial^{2}\mu_{i\ell}}{\partial\tilde{x}_{k}\partial\beta_{rs}}\cdot\sum_{j\neq i}[S(\rho)^{-1}]_{ij}.$
This uses the same $\frac{\partial^{2}\mu_{i\ell}}{\partial\tilde{x}_{k}\partial\beta_{rs}}$ as for direct effects.
The derivative with respect to $\rho$ is
$\frac{\partial IE_{i\ell k}}{\partial\rho}=\frac{\partial}{\partial\rho}\left[\frac{\partial\mu_{i\ell}}{\partial\tilde{x}_{k}}\right]\cdot\sum_{j\neq i}[S(\rho)^{-1}]_{ij}+\frac{\partial\mu_{i\ell}}{\partial\tilde{x}_{k}}\cdot\sum_{j\neq i}\frac{\partial[S(\rho)^{-1}]_{ij}}{\partial\rho},$
where
$\sum_{j\neq i}\frac{\partial[S(\rho)^{-1}]_{ij}}{\partial\rho}=\sum_{j\neq i}[S(\rho)^{-1}WS(\rho)^{-1}]_{ij}.$